### v3

# WritingBench: A Comprehensive Benchmark for Generative Writing

Abstract
Recent advancements in large language models (LLMs) have significantly enhanced text generation capabilities, yet evaluating their performance in generative writing remains a challenge. Existing benchmarks primarily focus on generic text generation or limited in writing tasks, failing to capture the diverse requirements of high-quality written contents across various domains. To bridge this gap, we present WritingBench, a comprehensive benchmark designed to evaluate LLMs across 6 core writing domains and 100 subdomains, encompassing creative, persuasive, informative, and technical writing. We further propose a query-dependent evaluation framework that empowers LLMs to dynamically generate instance-specific assessment criteria. This framework is complemented by a fine-tuned critic model for criteria-aware scoring, enabling evaluations in style, format and length. The framework’s validity is further demonstrated by its data curation capability, which enables 7B-parameter models to approach state-of-the-art (SOTA) performance. We open-source the benchmark, along with evaluation tools and modular framework components, to advance the development of LLMs in writing.
WritingBench: A Comprehensive Benchmark for Generative Writing
Yuning Wu1, Jiahao Mei1,3, Ming Yan1thanks: Corresponding authors, Chenliang Li1, Shaopeng Lai1, Yuran Ren2, Zijia Wang2, Ji Zhang1, Mengyue Wu3, Qin Jin2*, Fei Huang1 1Alibaba Group, 2Renmin University of China, 3Shanghai Jiao Tong University {yuningwu, ym119608}@alibaba-inc.com, qjin@ruc.edu.cn
1 Introduction
Figure 1: WritingBench query example with color-coded requirements. The three black-bordered categories highlight essential requirements analyzed in follow-up assessments. Red phrases correlate with gray-shaded writing support materials.
Figure 2: Construction pipeline of WritingBench queries. The refinement pool contains five writing requirements (three core competencies with black borders) and one expression type (purple). Checked strategies refine initial queries into multi-requirement prompts (color-coded text) with red phrases referencing gray materials. Implementation details in Section 3.1.
In recent years, LLMs DeepSeek-AI et al. (2025); Hurst et al. (2024) have revolutionized text generation, demonstrating impressive performance across diverse applications, from generating creative content Mirowski et al. (2023); Lei et al. (2024) and assisting in education Wang et al. (2024a); Li et al. (2024b) to enhancing professional workflows Shao et al. (2024); Li et al. (2024a). However, generative writing, which requires high levels of coherence, creativity, logical reasoning, and stylistic precision, poses a unique challenge that existing evaluation methods fail to address adequately.
Current evaluation benchmarks for generative writing suffer from two major limitations: 1) Limited scope and diversity in task formulation; and 2) Inadequate evaluation metrics for complex writing tasks. First, there is a significant lack of specialized benchmarks covering a broad range of writing tasks. Most existing writing-oriented benchmarks are restricted to single domains, such as fictions Karpinska et al. (2024); Gómez-Rodríguez and Williams (2023), and their task formulations tend to be simplistic—often relying on single-sentence queries Bai et al. (2024) or a small set of instruction templates Paech (2023); Que et al. (2024). Additionally, many benchmarks use homogeneous input materials Que et al. (2024); Karpinska et al. (2024), limiting their ability to accommodate the complex and customized requirements inherent in real-world writing scenarios. As a result, they fail to capture the diversity and intricacies of practical writing tasks (see Figure 1). Second, current automatic evaluation metrics lack the robustness needed for a comprehensive and nuanced assessment of writing quality. While LLM-based evaluation methods show promise in capturing semantic meanings Shao et al. (2024); Que et al. (2024); Bai et al. (2024), they typically rely on a narrow set of predefined criteria (e.g., fluency and coherence). As LLMs continue to evolve with increasingly sophisticated writing capabilities, these static evaluation criteria and frameworks are inadequate for assessing the complex, multi-dimensional nature of writing, including creativity, argumentation strength, and domain-specific adherence.
To address these challenges, we introduce WritingBench, a comprehensive benchmark and robust framework for evaluating general-purpose writing. Our approach begins with a carefully designed secondary domain categorization, grounded in real-world writing needs. We develop a four-stage query construction pipeline (illustrated in Figure 2), where LLMs first generate and diversify writing queries, followed by human-driven material collection and optimization. This process ensures a diverse set of writing tasks with broad domain coverage, varied requirements, and integration of heterogeneous source of materials. To enable a more nuanced evaluation, we propose a query-dependent evaluation framework that dynamically generates five instance-specific criteria using LLMs, which are then scored by a fine-tuned critic model. Finally, we integrate the framework to filter writing-specific data and train a small-scale model to verify its ability in identifying high-quality writing samples.
Our primary contributions are as follows:
$\bullet$ We present WritingBench, an open-source writing benchmark comprising 1,239 queries across 6 primary domains and 100 subdomains,featuring style, format and length requirements. WritingBench supports extended-context generation with input ranging from tens to thousands of words, addressing real-world diversity. It facilitates systematic evaluation to identify improvement areas and highlights the potential of chain-of-thought (CoT) processes in creative tasks.
$\bullet$ We propose a query-dependent evaluation framework that integrates instance-specific criteria generation with a criteria-aware scoring model. It achieves 83% human alignment, significantly surpassing static-criteria baselines (65%, 59%). The effectiveness is further evidenced by its data curation capability - models trained with framework-filtered data match SOTA performance.
$\bullet$ We publicly release WritingBench, including its evaluation protocols, criteria generation tools with an integrated critic model, and writing-enhanced models, to foster further research. Available at: https://github.com/X-PLUG/WritingBench.
2 Related Work
2.1 Writing Benchmarks
Existing evaluation benchmarks suffer from significant limitations in domain coverage and task granularity. For instance, EQ-Bench encompasses story-related quires using templated queries Paech (2023), while LongBench-Write incorporates length constraints in 120 queries Bai et al. (2024); however, they both lack hierarchical domain taxonomies and multi-dimensional requirement specifications (e.g., style and format). Furthermore, most current benchmarks rely on fixed instruction templates, short contexts or materials predominantly from a single source Que et al. (2024); Karpinska et al. (2024); Marco et al. (2024), rendering them insufficient for addressing the complexity of real-world data needs. In contrast, our proposed benchmark fills these gaps by introducing 1,239 free-form queries distributed across 6 primary domains and 100 subdomains, with potential controls over style, format, and length, paired with inputs ranging from tens to thousands of words.
Benchmark Num Domains Requirement Input Token Free Query-Form Diverse Material-Source
Primary Secondary Style Format Length Avg Max
EQ-Bench 241 1 / 130 213 /
LongBench-Write 120 7 / 87 684 /
HelloBench 647 5 38 1,210 7,766
WritingBench (Ours) 1,239 6 100 1,546 19,361
Table 1: Comparison of existing writing benchmarks.
2.2 Evaluation Methods
Using LLMs as evaluators has become a prevalent approach for evaluating the quality of generated responses. Typically, researchers pre-define a fixed set of evaluation dimensions applicable across all test instances Gómez-Rodríguez and Williams (2023); Marco et al. (2024). For example, SuperCLUE Xu et al. (2023) employs three dimensions, whereas LongBench-Write Bai et al. (2024) adopts six dimensions (e.g., relevance and accuracy). HelloBench Que et al. (2024) introduces task-specific dimensions, but the dimensions remain consistent across all queries of a given task. Although the LLM-as-a-judge approach enhances scalability, static evaluation dimensions often fail to accommodate the diversity of writing styles and specifications. To address this limitation, recent work Liang et al. (2024) trains a model to dynamically generate evaluation dimensions for individual queries. However, the dimensions remains confined to a small predefined set. In contrast, our query-dependent evaluation framework leverages LLMs to generate diverse and instance-specific criteria while fine-tuning a dedicated critic model to perform the evaluation.
2.3 Writing-Enhanced Models
Although existing LLMs demonstrate exceptional writing capabilities, researchers strive to further enhance their overall writing proficiency. Recent models, such as Weaver Wang et al. (2024b) and Suri Pham et al. (2024), have exhibited notable domain-specific strengths. For instance, Weaver benefits from over 200B parameter pretraining, supporting four distinct writing domains, while LongWriter Bai et al. (2024) specializes in length constraints. However, these models experience substantial performance degradation in cross-domain scenarios and multi-constraint tasks. In this work, leveraging the effectiveness of our evaluation framework, we introduce writing-enhanced models trained on high-quality data, achieving high performance across various tasks.
3 WritingBench
Figure 3: Domain categories in WritingBench. Inner represents the 6 primary domains and outer depicts 100 secondary subdomains (Sub = subdomains per category, Num = dataset entries).
Category Num
Avg
Token
Max
Token
Domain
Academic & Engineering 187 1,915 15,534
Finance & Business 238 1,762 19,361
Politics & Law 226 2,274 18,317
Literature & Arts 242 1,133 9,973
Education 151 1,173 10,737
Advertising & Marketing 195 886 6,504
Requirement
Style 395 1,404 18,197
Format 342 1,591 18,197
Length 214 1,226 14,097
Length
<1K 727 443 994
1K-3K 341 1,808 2,991
3K-5K 94 3,804 4,966
5K+ 77 8,042 19,361
Table 2: Data statistics for WritingBench categorized by domain, requirement, and length.
In this section, we will mainly introduce the construction process of our WritingBench, and the query-dependent evaluation framework with the accompanied critic model for criteria-aware evaluation. Additionally, we train a writing-enhanced model to demonstrate the effectiveness of our framework’s data curation capability.
3.1 Benchmark Construction
WritingBench is constructed following a delicate pipeline (see Figure 2) that combines model-generated query refinement and human annotation, ensuring diversity and real-world relevance. The two-phrase construction process is illustrated below.
3.1.1 Model-Augmented Query Generation
This phase leverages LLMs to generate an initial set of writing queries, which are enriched and diversified through systematic guidance, with material suggestions provided as needed.
Phase 1: Initial Query Generation
We begin by constructing a two-tiered domain pool grounded in real-world writing scenarios, consisting of 6 primary domains and 100 secondary subdomains (see Figure 3 and Appendix A for detailed descriptions). The selected domains are designed to capture both traditional and emerging user needs for AI-assisted writing, encompassing categories such as Academic & Engineering, Finance & Business, Politics & Law, Literature & Art, Education, Advertising & Marketing. Leveraging the domain and subdomain tags, we prompt two different LLMs (ChatGPT-4o-latest and Claude-3.5-Sonnet) to generate initial writing queries that simulate realistic user requests to maximize diversity (see Appendix B.1).
Phase 2: Query Diversification
To improve the diversity and practical applicability of queries, while taking into account real-world needs, we design a set of query diversification strategies (see Appendix B.2) inspired by Xu et al. (2024), include three core requirements and three auxiliary dimensions:
$\bullet$ Stylistic adjustments (e.g., “Use a friendly and simple tone that kids can easily understand.”)
$\bullet$ Format specifications (e.g., “Follow the IEEE conference template”)
$\bullet$ Length constraints (e.g., “Generate a 500-word executive summary”)
$\bullet$ Personalization (e.g., “From the viewpoint of an educator with two decades of experience.”)
$\bullet$ Content specificity (e.g., “Detail the 2023 Q3 financial metrics”)
$\bullet$ Expression (e.g., Modify the query expression to be shorter.)
The first three categories represent typical writing capabilities, which we will evaluate through specialized assessments. While refining the queries, the LLM simultaneously provides material requirements (e.g., requesting financial reports as input for market analysis queries) (see Appendix B.3). This approach results in enriched queries accompanied by corresponding recommended reference materials.
Figure 4: Example of dynamically generating criteria for a writing query in WritingBench. Different background colors represent various types of requirements.
3.1.2 Human-in-the-Loop Refinement
This phase incorporates human expertise to verify model-generated queries and supplement material requirements, thereby ensuring their alignment with real-world applications and avoiding harmful content.
Phase 1: Material Collection
At this stage, we engage 30 trained annotators, each compensated at \$18 per hour and possessing specialized expertise proven through domain-knowledge tests. Their role is to collect necessary open-source materials (e.g., public financial statements or legal templates), guided by material requirements generated by LLMs. To minimize errors arising from parsing documents in diverse formats, the annotators carefully extract and verify the most pertinent text segments.
Phase 2: Expert Screening & Optimization
Next, five experts with LLM experience or industry expertise perform data screening. The experts conduct a delicate two-stage filtering process: (1) query adaptation: ambiguous or unrealistic queries are revised to better align with the provided materials and practical scenarios (e.g., adjusting a legal opinion query to reference specific clauses from the supplied statutes), (2) material pruning: redundant or irrelevant content is eliminated from the collected materials, ensuring that the context provided for writing tasks remained focused and relevant while guaranteeing the information is free from harmful content.
Finally, we construct WritingBench, a benchmark comprising 1,239 queries categorized using a two-tiered taxonomy. In comparison to existing writing benchmarks summarized in Table 1, WritingBench exhibits notable advantages in terms of the number of instances, domain diversity, requirement coverage, and variability in input lengths. The detailed statistical distribution of WritingBench is shown in Table 2.
3.2 Query-Dependent Evaluation Framework
Traditional LLM-as-a-judge evaluations typically rely on fixed evaluation criteria derived from general writing assessment conventions Shao et al. (2024); Bai et al. (2024); Que et al. (2024). However, such static criteria exhibit three critical limitations: (1) Domain exhaustiveness: fixed criteria are inadequate in adapting to specialized domains, failing to address unique characteristics found in areas like technical documentation or creative writing; (2) Requirement specificity: such criteria lack the flexibility to encompass specific requirements related to style, format, or length control; and (3) Material dependency: they are insufficient to verify whether responses appropriately utilize the provided reference materials, such as incorporating data points or maintaining narrative continuity.
To address these challenges, we propose a query-dependent evaluation framework that enables dynamic adaptation to diverse writing scenarios. Our approach comprises two phases:
Phase 1: Dynamic Criteria Generation
As illustrated in Figure 4, given a query $q$ in the WritingBench, the LLM is prompted to automatically generate a set of five evaluation criteria, $C_{q}=\{c_{1},\ldots,c_{5}\}$ (opting for five criteria reflects a common approach observed in many contemporary evaluation contexts Shao et al. (2024); Ma et al. (2024); Marco et al. (2024)), using a carefully designed instruction to ensure structural guidance during criteria specification (see Appendix B.4). Each criterion comprises three components: a concise name summarizing the criterion, an extended description elaborating on the evaluation focus, and detailed scoring rubrics, which provide fine-grained quality levels for the respective evaluation dimensions.
Phase 2: Rubric-based Scoring
For each criterion $c_{i}\in C_{q}$, the evaluator independently assigns a score on a 10-point scale to a response $r$, providing both a score and a justification. The final overall score is computed by averaging scores across all criteria. Detailed prompt is provided in Appendix B.5.
Models Avg Languages Domains Requirements
ZH EN D1 D2 D3 D4 D5 D6 R1 C R2 C R3 C
Proprietary LLMs
ChatGPT-4o-latest 8.16 8.3 8.1 8.1 8.1 8.2 8.1 8.4 8.1 8.3 8.7 8.2 8.9 8.2 8.3
o1-Preview 8.15 8.1 8.2 8.0 8.1 8.2 8.2 8.4 8.1 8.2 8.6 8.2 8.8 8.2 8.2
Claude-3.5-Sonnet 7.71 7.7 7.7 7.6 7.5 7.6 7.7 7.9 8.0 7.9 8.5 7.7 8.5 7.9 8.0
Gemini-1.5-Pro 7.78 7.8 7.7 7.7 7.5 7.8 7.9 8.0 7.9 7.9 8.6 7.9 8.8 7.9 8.0
Qwen-Max 8.37 8.4 8.3 8.3 8.3 8.4 8.4 8.5 8.4 8.5 8.7 8.4 9.0 8.4 8.5
Open-source LLMs
Deepseek-R1 8.55 8.7 8.5 8.5 8.5 8.6 8.6 8.7 8.6 8.7 8.9 8.6 9.0 8.6 8.7
Deepseek-V3 7.95 8.0 7.9 7.9 7.8 8.0 7.8 8.2 8.0 8.1 8.6 8.0 8.9 8.0 8.2
Mistral-Large-Instruct 7.64 7.6 7.7 7.7 7.6 7.8 7.3 7.9 7.6 7.7 8.2 7.7 8.7 7.7 7.9
Qwen-2.5-72B-Instruct 7.90 8.0 7.9 8.0 7.8 8.1 7.7 8.2 7.8 8.0 8.3 8.0 8.8 7.9 8.0
Qwen-2.5-7B-Instruct 7.43 7.3 7.5 7.7 7.4 7.6 6.9 7.8 7.3 7.5 7.9 7.6 8.6 7.4 7.5
Llama-3.3-70B-Instruct 7.01 6.7 7.3 7.0 6.9 7.0 6.8 7.3 7.3 7.1 7.8 7.1 8.2 7.0 7.2
Llama-3.1-8B-Instruct 6.35 5.7 6.9 6.6 6.4 6.1 6.0 6.7 6.6 6.4 7.0 6.4 7.6 6.3 6.4
Capability-enhanced LLMs
Suri 4.97 4.4 5.5 5.6 5.3 5.0 4.1 5.0 5.1 4.8 5.2 5.0 5.4 4.5 4.0
LongWriter 7.91 7.9 7.9 8.0 8.1 8.1 7.7 8.1 7.6 7.9 8.2 8.1 8.8 7.7 7.7
Qwen-2.5-7B-filtered 8.49 8.6 8.4 8.4 8.4 8.6 8.4 8.6 8.5 8.6 8.8 8.5 9.0 8.5 8.6
Llama-3.1-8B-filtered 8.49 8.6 8.4 8.5 8.4 8.6 8.4 8.6 8.5 8.6 8.8 8.5 8.9 8.5 8.5
Table 3: WritingBench performance of different LLMs across 6 domains and 3 writing requirements evaluated with our critic model (scale: 1-10). The six domains include: (D1) Academic & Engineering, (D2) Finance & Business, (D3) Politics & Law, (D4) Literature & Art, (D5) Education, and (D6) Advertising & Marketing. The three writing requirements assessed are: (R1) Style, (R2) Format, and (R3) Length. Here, “C" indicates category-specific scores.
3.3 Critic Model
To alleviate the computational overhead with LLM-based evaluation, we develop a dedicated critic model, $\mathcal{M}$, designed to implement our rubric-based scoring framework. Specifically, this model performs the mapping $\mathcal{M}_{c}:(q,r,C_{i})\mapsto[1,10]\times\mathcal{J}$, where the output consists of a numerical score and corresponding justification text, $\mathcal{J}$, both in accordance with the predefined evaluation rubric.
We fine-tune the critic model on a dataset comprising 50K instances, which are collected using LLMs in our experiments. The dataset encompasses diverse queries, evaluation criteria, and model responses to enhance the robustness of evaluation. The experiments presented in Section 4.3 validate the consistency of the critic model.
3.4 Evaluation-Guided Data Curation for Writing Enhancement
The query-dependent evaluation framework enables systematic data curation across diverse writing tasks through two-phase filtering. First, applying instance-specific criteria generation to initial 24K SFT samples. Subsequently, the critic model filters 50% samples, yielding 12K framework-curated samples with higher quality. Fine-tuning experiments are conducted using the llama-3.1-8b-instruct and qwen-2.5-7b-instruct models. Evaluations are performed on two writing benchmarks to validate the framework’s effectiveness in identifying high-quality writing samples.
4 Experiment
Figure 5: Scores of different models across different subdomains in WritingBench. Red indicates higher score and blue refers to lower score. The figures are the average score of each subdomain for different models.
4.1 Experiment Settings
This section outlines the settings employed in our experiments using the WritingBench framework. The evaluation protocol, and the model training configurations are as follows:
Evaluation Protocol: Uniform generation settings are applied to model responses on WritingBench, allowing up to 16,000 tokens. We set the temperature at 0.7, with a top-k of 20 and top-p of 0.8, ensuring a balance of creativity. The critic model uses the same parameters, with input length capped at 2,048 tokens for scoring stability.
Model Training: The critic model is fine-tuned on the Qwen-2.5-7B-Instruct base model, using the AdamW optimizer with a 7e-6 learning rate. It is trained on 50K SFT data for 3 epochs across 8xA100 GPUs (batch size 64, 8-step accumulation). The writing models, trained on both Qwen-2.5-7B-Instruct and Llama-3.1-8B-Instruct for 5 epochs on 32xA100 GPUs, achieving a total batch size of 128 with 4-step gradient accumulation.
4.2 Comparison between LLMs
We evaluate 16 LLMs: ChatGPT-4o-latest Hurst et al. (2024), o1-Preview Jaech et al. (2024), Claude-3.5-Sonnet Anthropic (2024), Gemini-1.5-Pro Reid et al. (2024), Qwen-Max Team (2025), Deepseek-R1 DeepSeek-AI et al. (2025), Deepseek-V3 DeepSeek-AI et al. (2024), Mistral-Large-Instruct MistralAI (2024), Qwen-2.5-72B-Instruct and Qwen-2.5-7B-Instruct Yang et al. (2024), Llama-3.3-70B-Instruct and Llama-3.1-8B-Instruct Dubey et al. (2024), Suri Pham et al. (2024), LongWriter Bai et al. (2024), and our writing model, Qwen-2.5-7B fine-tuned model and Llama-3.1-8B fine-tuned model, on WritingBench. Each query is assessed by 5 instance-specific criteria using a 10-point scale, with subdomain-specific subcategory heatmaps revealing detailed variations.
Key Insights from Domain Scores: From Table 3, we observe that Education (D5) always yields high performance among models, followed by Academic & Engineering (D1). Literature & Art (D5) is the lowest-scoring domain, showing notable variance. CoT-capable models like Deepseek-R1 and o1-preview outperform their non-CoT counterparts models, indicating CoT’s potential in handling narrative and creative content. Figure 5 identifies challenges in niche areas such as novel continuations, tender proposals, and white papers where models generally score lower. These tasks demand a higher level of knowledge, long-text generation capabilities, and adherence to contextual consistency, pinpointing areas for further enhancement.
Key Insights from Requirement Scores: Table 3 presents scores for three requirement subsets: R columns showing overall scores and C columns showing requirement-specific averages. For example, if a style query includes two relevant criteria (as confirmed by human annotation), the C column averages scores for those criteria only. Style (R1) and Format (R2) dimensions reveal model distinctions , with Deepseek-R1 and Qwen-Max often excels others. Nevertheless, length requirements (C of R3) remains challenging, especially in section-specific constraints (see Appendix 16 for examples) and extended text production (see Section 4.5).
The overall analysis of WritingBench experiment highlights: Deepseek-R1 consistently leads across both domain and requirement dimensions, showcasing its versatility and strong language capabilities. There is significant performance variance within creative content domains, with CoT models showing notable promise—a hypothesis further confirmed by our comparison (see Appendix C.1, where the model incorporating CoT mechanisms outperform the one without CoT on creative writing tasks). Additionally, we observe that advanced models often achieve higher specialized scores (C column) for the three common requirements than their overall scores (R column). The criteria outside these specialized sets tend to emphasize content aspects, such as integration with source materials and writing depth, highlighting the need to improve content quality. This analysis benchmarks existing model capabilities and underscores specific areas needing improvement for future development.
4.3 Human Consistency
To validate the alignment between automated evaluation and human judgment, we conducted human evaluation on 300 queries, covering all 100 subdomains. Five professionally trained annotators with linguistic backgrounds perform pairwise comparisons of model responses. For each query, two responses are randomly selected from different models. Annotators select the preferred response or declared equivalence based on the query’s requirements. We compare our dynamic query-dependent criteria against two baselines: static globally uniform criteria and static domain-specific customized criteria (static criteria are designed by domain experts) on two LLM judges: ChatGPT-4o-latest and Claude-3.5-Sonnet.
Evaluation Metric Judge Score
Static Global ChatGPT 69%
Static Domain-Specific ChatGPT 40%
Dynamic Query-Dependent ChatGPT 79%
Static Global Claude 65%
Static Domain-Specific Claude 59%
Dynamic Query-Dependent Claude 87%
Dynamic Query-Dependent Critic Model 83%
Table 4: Comparison of human consistency scores across different criteria generation methods. ChatGPT corresponds to ChatGPT-4o-latest, Claude corresponds to Claude-3.5-Sonnet.
As shown in Table 4, our dynamic query-dependent criteria achieve superior human alignment compared to static, both globally uniform or domain-specific customized criteria. Notably, domain-specific criteria underperform despite customization, since our queries involve highly diverse tasks and varied sources. These findings confirm that context-sensitive query-dependent evaluation better captures real-world writing complexity compared to conventional static approaches. Furthermore, the critic model attains 83% agreement, confirming its practical viability.
4.4 Ablation of Data Curation for Writing-Enhanced Models
Leveraging the query-dependent evaluation framework described in Section 3.2, we perform instance-specific criteria generation an initial 24K SFT dataset spanning various domains, including long-input queries enriched with supplemental materials. Our critic model then apply a query-dependent evaluation to filter this dataset down to a high-quality 12K subset. We conduct fine-tuning experiments using the Llama-3.1-8B-Instruct and Qwen-2.5-7B-Instruct models. As shown in 5, both models, trained on the filtered data, demonstrate significant performance improvements over their previous versions. Notably, they outperform models trained on the full dataset and even approached the capabilities of SOTA models. Further validation on an alternative benchmark, LongBench-Write (following the settings of quality evaluation in Bai et al. (2024)), confirm these findings, with our models surpassing baseline performance.
These results validate the robustness of our query-dependent evaluation strategy and the utility of our critic model for curating high-quality writing samples. This integrated approach enables smaller models to compete with, and even surpass, larger models across diverse writing tasks.
Models WritingBench Benchmark2
Deepseek-R1 8.55 4.79
Qwen-2.5-7B-Instruct 7.43 4.39
Llama-3.1-8B-Instruct 6.35 3.12
Qwen-2.5-7B-all 8.46 4.69
Qwen-2.5-7B-filtered 8.49 4.70
Llama-3.1-8B-all 8.45 4.65
Llama-3.1-8B-filtered 8.49 4.65
Table 5: Performance evaluation of our writing models on two benchmarks.’-filtered’ indicates models trained with filtered data, while ’-all’ refers to those trained with the full dataset.
4.5 Ablation of Length
We assess the performance of LLMs across varying input and output lengths, with statistical validity ensured by excluding intervals containing fewer than 5 samples.As demonstrated in Appendix 6, experiments reveal that most SOTA models generally maintain consistent performance regardless of input length variations, attributable to their advanced long-context comprehension capabilities.
However, analysis in Appendix 7 shows that most models exhibit inherent limitations in response generation length, typically producing outputs constrained to approximately 3,000 tokens. While output quality remains relatively stable within this range, smaller models suffer more performance degradation characterized by repetitive outputs. Notably, only two models, Qwen-Max and LongWriter, which incorporate architectural optimizations for long-form generation, effectively support extended response lengths. It underscores the importance of enhancing long-output capabilities and length optimization in writing.
5 Conclusion
In this paper, we introduce WritingBench, a comprehensive benchmark designed to evaluate LLMs in generative writing across diverse domains. It includes 1,239 queries spanning 6 primary domains and 100 subdomains, providing evaluation dimensions for style, format, and length requirements. Our query-dependent evaluation framework, supported by a critic model, achieves high human alignment. Evaluation efficiency is further demonstrated by compact models trained on curated data, achieving performance comparable to SOTA models. By making WritingBench and its resources publicly available, we aim to foster further research and advancements in LLM writing capabilities.
Limitations
This work faces three primary limitations that warrant consideration. First, both the writing model and critic model are primarily trained using conventional SFT approaches, omitting systematic exploration of enhanced optimization strategies. While we demonstrate the partial efficacy of the CoT mechanisms in creative domains, their potential remains unexplored compared to established successes in mathematical reasoning tasks.
Second, our evaluation framework exhibits diminished precision in handling complex multi-dimensional length requirements, including temporal sequencing constraints and section-specific word counts. This limitation underscores the necessity for enhanced scoring methodologies that integrate learned metrics with structured rule-based evaluations to better regulate output specifications.
Third, inherent challenges persist in obtaining reliable pairwise preference annotations for compositional tasks. Despite rigorous annotation protocols, human evaluators inevitably introduce subjective biases when assessing two fair-well responses, particularly regarding narrative preferences and contextual interpretations. While our consensus-building procedures mitigate some variability, absolute alignment with diverse user preferences remains theoretically unattainable.
Appendix A Benchmark Statistics
A.1 Overview of Six Primary Domains
  1. 1.
    Academic & Engineering: This domain encompasses the structured and formalized nature of academic writing workflows, focusing on clarity, precision, and adherence to rigorous standards. includes the creation of paper outlines, abstracts, literature reviews, experiment reports, and technical documents such as patents and test reports. The writing prioritizes logical argumentation, thorough analysis, and the integration of empirical evidence.
  2. 2.
    Finance & Business: Writing in this domain is analytical and strategic, aimed at informing decision-making and promoting corporate objectives. It includes a wide range of documentation such as contracts, market analyses, investment reports, strategic plans, and operational materials like product specifications and sales reports. The emphasis is on clarity and conciseness, with a focus on financial acumen and strategic insights.
  3. 3.
    Politics & Law: This domain demands an authoritative and formal tone, as it involves the composition of government documents, legal writings, and political communications. These materials require a careful balance between clarity and formality, often employing complex and structured language. The aim is to clearly convey policy positions, legal arguments, or political messages while strictly adhering to legal and procedural standards.
  4. 4.
    Literature & Art: This domain covers the creative and expressive realms of writing, including novels, poetry, scripts, artistic designs, and critiques of books and movies. Writers explore thematic and emotional depths, crafting works that connect with audiences on a human level. The language is rich and evocative, allowing for a personal exploration of ideas that engage and move the reader.
  5. 5.
    Education: This domain involves pedagogical materials and educational communication, including lesson plans, course designs, feedback, assignments, and institutional communications like admissions promotions and parent-teacher meeting scripts. The writing prioritizes clarity, accessibility, and instructional effectiveness, using an approachable tone to facilitate learning and engagement.
  6. 6.
    Advertising & Marketing: Writing in this domain is vibrant and persuasive, designed to captivate and influence target audiences across various digital platforms. It includes social media scripts, advertising copy, brand narratives, and multimedia campaign materials. The writing is dynamic and strategic, with a creative twist, necessitating a deep understanding of audience psychology and trend dynamics to ensure that the content is appealing and strategically effective.
A.2 Overview of 100 Secondary Subdomains
See Table 7, Table 8 and Table 9.
Appendix B Prompts
B.1 Initial Query Generation Prompt
Initial Query Generation Prompt Generate 10 different writing requests under {domain2} within the context of {domain1}. Ensure the requests are as detailed and specific as possible, and reflect realistic user tone and needs.
Please return in the following JSON format, and do not include anything outside of JSON:
   "Writing request 1",
   "Writing request 2",
Table 6: Initial query generation prompt introduced in Section 3.1.1.
Subdomain Description
Academic & Engineering
Paper Outline Hierarchical organization of research components and logical flow
Acknowledgments Formal recognition of institutional and individual support
Limitations Systematic identification of methodological constraints and scope boundaries
Defense Presentation Presentation supporting materials, such as slides
Research Proposal Investigation blueprint with validation road map
Technical Documentation Implementation specifications and system interface protocols
Experiments Parameterized validation framework with controlled variable analysis
Introduction Contextual foundation establishing research gaps and significance
Conclusion synthesize the main findings of the research or project
Test Report Evaluations of testing activities and performance
Contributions Novel aspects differentiating the work from prior research
Internship Report Chronological documentation of a practical work placement
Literature Review Critical gap analysis through scholarly works taxonomy
Defense Script Oral presentations and responses for research defense.
Abstract Summary of research objectives, methods, results, and significance
Engineering Report Technical analysis on tasks, methodologies, and outcomes
Patent Legal-technical specification of novel implementable claims
Finance & Business
Meeting Minutes Concise documentation of key discussion points, decisions, and action items
User Research Insight collection on user needs and behaviors to inform product or service design
Business Correspondence Formal communication with internal or external stakeholders for business purposes
Human Resource Management Strategies and processes for managing workforce effectively
Recruitment Strategies for attracting, selecting, and onboarding suitable candidates
Briefing Summarized information provided to stakeholders ahead of a task or meeting
Event Planning Coordinated organization of logistics and activities for event execution
Market Research Systematic collection and analysis about market and consumer
Market Analysis Evaluation of market trends, size, competitors, and dynamics
Risk Management Identification, assessment, and prioritization of risks with mitigation strategies
Sales Report Summary of sales activities, performance, and revenue figures over a given period
Pitch Deck Visual presentation designed to communicate business ideas or proposals to investors
Contract Legally binding agreement outlining the terms and conditions for business transactions
Tender Document Formal proposal request containing project specifications and bidding instructions
Investment Analysis Evaluation of financial investments to determine potential returns and risks
Product Proposal Detailed plan outlining the development, features, and potential of new products
Strategic Planning Business goal setting with actionable strategies for desired outcomes
Financial Reports Comprehensive statements reflecting the financial performance and status
Requirements Specification Documentation detailing functional and non-functional requirements for a project
Bid Proposal Formal offer to provide goods or services at a specified price, addressing client needs
Table 7: Subdomains in Academic & Engineering and Finance & Business.
Domain Description
Politics & Law
Legal Opinion Authoritative assessment and guidance on legal matters or questions
Government Speech Formal address delivered by government officials outlining policies or positions
Judgment Document Official written decision or order issued by a court
Legal Agreement Binding contract setting out terms and obligations between parties
Case Study In-depth analysis of a legal case for educational or professional purposes
Case Bulletin Summary and update on ongoing or concluded legal cases
Legal Consultation Professional advice provided on legal rights, responsibilities, or strategies
Regulatory Analysis Examination of rules and regulations affecting compliance and enforcement
Meeting Summary Brief overview of discussions, decisions, and outcomes from a meeting
Ideological Report Analysis or commentary on political or ideological trends and perspectives
Policy Interpretation Explanation or clarification for public or organizational guidance
Official Document Formal written record issued by government entities or officials
Legal Awareness Campaign Initiative to educate the public on legal rights and responsibilities
Defense Plea Formal written argument submitted by the defense in a legal proceeding
Party Membership Application Form and process for joining a political party
Policy Advocacy Efforts to influence or promote specific policy changes or implementations
Work Report Detailed account of activities, achievements, and challenges within a specific period
Deed Achievement Record highlighting significant accomplishments and contributions
Litigation Documents Legal filings and paperwork submitted in the course of a lawsuit
White Paper Authoritative report providing information or proposals on an issue
Literature & Art
Character Design Creation and development of detailed characters for stories or visual media
Greeting Message Friendly or formal introductory statement used for various occasions
Host Script Guided narration and dialogue for a presenter during an event or show
Novel Outline Structured plan for the plot, characters, and settings of a novel
Podcast Script Written content outlining the dialogue and segments for podcast episodes
Derivative Work Creative work based on or inspired by an existing piece
Reading Reflection Personal thoughts and analysis on a piece of literature
Video Script Script detailing dialogue and action for video content creation
Book Review Critical evaluation and summary of a book’s content and impact
Game Design Creation of game mechanics, stories, and interfaces for interactive entertainment
Lyric Writing Crafting of words for songs with rhyme and meter considerations
Brainstorm Rough ideas and notes generated during a creative thinking session
Plot Development Process of mapping out the storyline and narrative structure
Prose Written or spoken language in its ordinary form, without metrical structure
Screenplay Scripted blueprint for film or television with dialogue and directions
Novel Manuscript Complete text of a novel prepared for publication
Biography Detailed account of a person’s life experiences and achievements
Film/TV Review Analytical critique of a film or television show’s content and effectiveness
Poetry Artistic composition using rhythmic and metaphorical language
Fan Fiction Amateur stories written by enthusiasts featuring characters from existing media
Table 8: Subdomains in Politics & Law and Literature & Art.
Domain Description
Education
Training Reflection Personal assessment of training experiences and learned insights
Class Activity Planned exercises or tasks designed to engage students in learning
Parent-Teacher Meeting Formal discussion between educators and parents about student progress
Lesson Plan Structured outline of educational objectives and teaching methods for a class
Teaching Materials Resources used to aid in presenting information to students
Assignment Grading Evaluation and scoring of student work based on specific criteria
Curriculum Design Development of educational content, structure, and delivery methods
Educational Report Analysis or summary of educational outcomes and performance
Coursework Academic work assigned to students as part of a course
Evaluation Comments Feedback provided on student performance and areas of improvement
Educational Consulting Professional guidance on educational strategies and systems
Admissions Promotion Strategies and activities aimed at encouraging enrollment in educational institutions
Advertising & Marketing
Sales Letter Persuasive written communication intended to motivate potential buyers
Product Description Detailed overview of a product’s features, benefits, and uses
Social Media Content Engaging text, images, or videos crafted for online platforms
Multimedia Script Planned screenplay integrating various forms of media for marketing
Promotional Copy Compelling text written to boost interest and sales of products
Promotional Voiceover Recorded narration to accompany marketing visuals or ads
Travel Guide Informative content offering insights and tips for travelers
Brand Story Narrative that outlines the history, values, and mission of a brand
Personal Blog Individual commentary or stories shared in an informal online format
Marketing Commentary Analytical thoughts on marketing trends and strategies
Slogans Catchy and memorable phrases designed to convey brand identity
Table 9: Subdomains in Education and Advertising & Marketing.
B.2 Guidance Pool
Query Refinement Guidance Pool [
    "Add a requirement for generating specific lengths.",
    "Include format adherence requirements, such as writing according to a prescribed outline or outputting in a specific format.",
    "Add style requirements, like drafting a speech suitable for a particular occasion or adopting the style suitable for a specific audience or mimicking a particular tone.",
    "Incorporate user personalization needs, such as considering the user’s identity or integrating personal experiences.",
    "Include more specific content requirements, like details about a particular event or focusing on specific content.",
    "Express concisely in one sentence."
Table 10: Query refinement guidance pool introduced in Section B.5.
B.3 Query Refine Prompt
Query Refinement Prompt Please refine and enhance the original writing requirements in the context of generating content in {domain2} from {domain1} based on the provided guidance. Include as many details as possible and indicate whether additional writing materials are needed.
** Original Writing Requirements **
{query}
** Guidance for Modification **
{guidance}
** Output Requirements **
Return the result strictly in the following JSON format, with no additional content outside the JSON:
{     "query": "Modified writing requirements",
    "material": "Whether additional reference materials are needed to supplement the writing requirements. If needed, provide suggestions for the materials; if not needed, return"
Table 11: Query refinement prompt introduced in Section 3.1.1.
B.4 Criteria Generation Prompt
System Evaluation Prompt You are an expert evaluator with extensive experience in evaluating the response of a given query.
Table 12: Evaluation system prompt introduced in Section 3.2.
Criteria Generation Prompt Please generate five strict evaluation criteria for assessing the response given the following query. Each criterion should include the following fields: name, criteria_description, 1-2, 3-4, 5-6, 7-8, 9-10.
The criteria should be designed to emphasize detailed assessment and distinguish subtle differences in quality. Ensure that the criteria can discern issues such as relevance, coherence, depth, specificity, and adherence to the query context.
Do not include any additional text. Only output the criteria in the specified JSON format.
** Query **
{query}
** Output format **
   {"name": "first_criteria_name",
    "criteria_description": "Description for the first criteria, emphasizing detailed and critical assessment.",
    "1-2": "Low score description: Clearly deficient in this aspect, with significant issues.",
    "3-4": "Below average score description: Lacking in several important areas, with noticeable problems.",
    "5-6": "Average score description: Adequate but not exemplary, meets basic expectations with some minor issues.",
    "7-8": "Above average score description: Generally strong but with minor shortcomings.",
    "9-10": "High score description: Outstanding in this aspect, with no noticeable issues." },
Table 13: Criteria generation prompt introduced in Section 3.2.
B.5 Rubric-based Scoring Prompt
Scoring Prompt Evaluate the Response based on the Query and criteria provided.
** Criteria **
{criteria}
** Query **
{query}
** Response **
{response}
Provide your evaluation based on the criteria:
{criteria}
Provide reasons for each score, indicating where and why any strengths or deficiencies occur within the Response. Reference specific passages or elements from the text to support your justification.
Ensure that each reason is concrete, with explicit references to the text that align with the criteria requirements.
Scoring Range: Assign an integer score between 1 to 10
Output format:
Return the results in the following JSON format, Only output this JSON format and nothing else:
    "score": an integer score between 1 to 10,
    "reason": "Specific and detailed justification for the score using text elements."
Table 14: Scoring prompt introduced in Section 3.2.
Appendix C Experiment Results
C.1 Ablation of CoT in Creative content
To validate the impact of CoT reasoning on creative content generation, we conduct knowledge distillation experiments using DeepSeek-R1. Two variants of Qwen-2.5-32B-Instruct are developed: 1) Base model trained with CoT-formatted instructions, 2) ablated version (-w/o CoT) using direct response generation. These models are evaluated on the Literature & Art (D5) subset of WritingBench and EQBench, a specialized benchmark for fiction evaluation, to thoroughly examine their capabilities in generating creative content.
Models WritingBench-D4 EQBench
Deepseek-R1 8.55 84.99
Qwen-2.5-32B-Instruct 7.34 48.17
Qwen-2.5-32B-CoT 8.66 82.48
-w/o CoT 8.49 79.43
Table 15: Ablation of CoT in creative content on two benchmarks.
As shown in Table 15, the CoT-enhanced model achieves 8.66 (vs. 8.49) on WritingBench-D5 and 82.48 (vs. 79.43) on EQBench, showing CoT’s effectiveness in generating creative content.
C.2 Case Study
Query Create advertising slogans for a pet grooming salon, including: - Main slogan (10-15 words) - Sub-slogan (15-20 words) - Service tagline (8-12 words) - Brand promise (10-15 words) The slogans should reflect professional pet grooming and care services while emphasizing thoughtful care for pet owners. The slogans should be catchy, memorable, with warm and friendly wording, and a light, cheerful tone. Response **Main Slogan:** Pamper Your Pet, Delight Your Heart. *Sub-Slogan:** Expert grooming and loving care, making your furry friend feel fabulous, inside and out. *Service Tagline:** Gentle paws, happy pets, stress-free smiles. *Brand Promise:** Quality grooming, compassionate care, every time.
Table 16: Bad case of length requirement. Query from WritingBench with index 539, response generated by Gemini-1.5-Pro.
Figure 6: Scores of different models across various input lengths on the WritingBench.
Figure 7: Scores of different models across various output lengths on the WritingBench.

### v4

@@
 Existing benchmarks primarily focus on generic text generation or limited in writing tasks, failing to capture the diverse requirements of high-quality written contents across various domains.
-To bridge this gap, we present WritingBench, a comprehensive benchmark designed to evaluate LLMs across 6 core writing domains and 100 subdomains, encompassing creative, persuasive, informative, and technical writing.
-We further propose a query-dependent evaluation framework that empowers LLMs to dynamically generate instance-specific assessment criteria.
+To bridge this gap, we present WritingBench, a comprehensive benchmark designed to evaluate LLMs across 6 core writing domains and 100 subdomains.We further propose a query-dependent evaluation framework that empowers LLMs to dynamically generate instance-specific assessment criteria.
 This framework is complemented by a fine-tuned critic model for criteria-aware scoring, enabling evaluations in style, format and length.
-The framework’s validity is further demonstrated by its data curation capability, which enables 7B-parameter models to approach state-of-the-art (SOTA) performance.
+The framework’s validity is further demonstrated by its data curation capability, which enables a 7B-parameter model to outperform the performance of GPT-4o in writing.
 We open-source the benchmark, along with evaluation tools and modular framework components, to advance the development of LLMs in writing.
-WritingBench: A Comprehensive Benchmark for Generative Writing
-Yuning Wu1, Jiahao Mei1,3, Ming Yan1thanks: Corresponding authors, Chenliang Li1, Shaopeng Lai1, Yuran Ren2, Zijia Wang2, Ji Zhang1, Mengyue Wu3, Qin Jin2*, Fei Huang1 1Alibaba Group, 2Renmin University of China, 3Shanghai Jiao Tong University {yuningwu, ym119608}@alibaba-inc.com, qjin@ruc.edu.cn
+GitHub Repo: https://github.com/X-PLUG/WritingBench
+Leaderboard: https://huggingface.co/spaces/WritingBench/WritingBench
 1 Introduction
-Figure 1: WritingBench query example with color-coded requirements.
-The three black-bordered categories highlight essential requirements analyzed in follow-up assessments.
-Red phrases correlate with gray-shaded writing support materials.
+Figure 1: Query example with color-coded requirements. Red phrases correlate with gray-shaded writing support materials.
+In recent years, LLMs (r1; 4ol) have revolutionized text generation, demonstrating impressive performance across diverse applications, from generating creative content (theatre-human; ex3) and assisting in education (education-survey1; education-survey2) to enhancing professional workflows (storm; metareview).
+However, generative writing, which requires high levels of coherence, creativity, logical reasoning, and stylistic precision, poses a unique challenge that existing evaluation methods fail to address adequately.
 Figure 2: Construction pipeline of WritingBench queries.
@@
 Implementation details in Section 3.1.
-In recent years, LLMs DeepSeek-AI et al.
-(2025); Hurst et al.
-(2024) have revolutionized text generation, demonstrating impressive performance across diverse applications, from generating creative content Mirowski et al.
-(2023); Lei et al.
-(2024) and assisting in education Wang et al.
-(2024a); Li et al.
-(2024b) to enhancing professional workflows Shao et al.
-(2024); Li et al.
-(2024a).
-However, generative writing, which requires high levels of coherence, creativity, logical reasoning, and stylistic precision, poses a unique challenge that existing evaluation methods fail to address adequately.
 Current evaluation benchmarks for generative writing suffer from two major limitations: 1) Limited scope and diversity in task formulation; and 2) Inadequate evaluation metrics for complex writing tasks.
-First, there is a significant lack of specialized benchmarks covering a broad range of writing tasks.
-Most existing writing-oriented benchmarks are restricted to single domains, such as fictions Karpinska et al.
-(2024); Gómez-Rodríguez and Williams (2023), and their task formulations tend to be simplistic—often relying on single-sentence queries Bai et al.
-(2024) or a small set of instruction templates Paech (2023); Que et al.
-(2024).
-Additionally, many benchmarks use homogeneous input materials Que et al.
-(2024); Karpinska et al.
-(2024), limiting their ability to accommodate the complex and customized requirements inherent in real-world writing scenarios.
+First, there is a significant lack of specialized benchmarks that cover a broad range of writing scenarios.
+Most existing writing-oriented benchmarks are restricted to single domains, such as fictions (nocha; confederacy-human), and their task formulations tend to be simplistic—often relying on single-sentence prompts (longwriter) or a small set of instruction templates (eqbench; hellobench).
+Additionally, many benchmarks use homogeneous input materials (hellobench; nocha), limiting their ability to accommodate the complex and customized requirements inherent in real-world writing.
 As a result, they fail to capture the diversity and intricacies of practical writing tasks (see Figure 1).
 Second, current automatic evaluation metrics lack the robustness needed for a comprehensive and nuanced assessment of writing quality.
-While LLM-based evaluation methods show promise in capturing semantic meanings Shao et al.
-(2024); Que et al.
-(2024); Bai et al.
-(2024), they typically rely on a narrow set of predefined criteria (e.g., fluency and coherence).
+While LLM-based evaluation methods show promise in capturing semantic meanings (storm; hellobench; longwriter), they typically rely on a narrow set of predefined criteria (e.g., fluency and coherence).
 As LLMs continue to evolve with increasingly sophisticated writing capabilities, these static evaluation criteria and frameworks are inadequate for assessing the complex, multi-dimensional nature of writing, including creativity, argumentation strength, and domain-specific adherence.
@@
 This process ensures a diverse set of writing tasks with broad domain coverage, varied requirements, and integration of heterogeneous source of materials.
-To enable a more nuanced evaluation, we propose a query-dependent evaluation framework that dynamically generates five instance-specific criteria using LLMs, which are then scored by a fine-tuned critic model.
+To enable a more nuanced evaluation, we propose a query-dependent evaluation framework where LLMs dynamically generates five instance-specific criteria, which are then scored by a fine-tuned critic model.
 Finally, we integrate the framework to filter writing-specific data and train a small-scale model to verify its ability in identifying high-quality writing samples.
 Our primary contributions are as follows:
-$\bullet$ We present WritingBench, an open-source writing benchmark comprising 1,239 queries across 6 primary domains and 100 subdomains,featuring style, format and length requirements.
+  • We present WritingBench, an open-source writing benchmark comprising 1,000 queries across 6 primary domains and 100 subdomains, featuring style, format and length requirements.
 WritingBench supports extended-context generation with input ranging from tens to thousands of words, addressing real-world diversity.
 It facilitates systematic evaluation to identify improvement areas and highlights the potential of chain-of-thought (CoT) processes in creative tasks.
-$\bullet$ We propose a query-dependent evaluation framework that integrates instance-specific criteria generation with a criteria-aware scoring model.
-It achieves 83% human alignment, significantly surpassing static-criteria baselines (65%, 59%).
-The effectiveness is further evidenced by its data curation capability - models trained with framework-filtered data match SOTA performance.
-$\bullet$ We publicly release WritingBench, including its evaluation protocols, criteria generation tools with an integrated critic model, and writing-enhanced models, to foster further research.
+  • We propose a query-dependent evaluation framework that integrates instance-specific criteria generation with a criteria-aware scoring model.
+It achieves 84% human alignment, significantly surpassing static-criteria baselines (67%, 58%).
+The effectiveness is further evidenced by its data curation capability-models trained with framework-filtered data outperform GPT-4o in writing.
+  • We publicly release WritingBench, including its evaluation protocols, criteria generation tools with an integrated critic model, and writing-enhanced models, to foster further research.
 Available at: https://github.com/X-PLUG/WritingBench.
@@
 2.1 Writing Benchmarks
-Existing evaluation benchmarks suffer from significant limitations in domain coverage and task granularity.
-For instance, EQ-Bench encompasses story-related quires using templated queries Paech (2023), while LongBench-Write incorporates length constraints in 120 queries Bai et al.
-(2024); however, they both lack hierarchical domain taxonomies and multi-dimensional requirement specifications (e.g., style and format).
-Furthermore, most current benchmarks rely on fixed instruction templates, short contexts or materials predominantly from a single source Que et al.
-(2024); Karpinska et al.
-(2024); Marco et al.
-(2024), rendering them insufficient for addressing the complexity of real-world data needs.
-In contrast, our proposed benchmark fills these gaps by introducing 1,239 free-form queries distributed across 6 primary domains and 100 subdomains, with potential controls over style, format, and length, paired with inputs ranging from tens to thousands of words.
+Existing writing benchmarks suffer from significant limitations in domain coverage and task granularity.
+For instance, EQ-Bench (referring to its creative writing subset) encompasses templated queries for story-related tasks (eqbench), while LongBench-Write incorporates length constraints in 120 queries (longwriter); however, they both lack hierarchical domain taxonomies and multi-dimensional requirement specifications (e.g., style and format).
+Furthermore, most benchmarks rely on fixed instruction templates, short contexts, or materials predominantly from a single source (hellobench; nocha; pros-human-5), rendering them insufficient for addressing the complexity of real-world needs.
+In contrast, our proposed benchmark fills these gaps by introducing 1,000 free-form queries distributed across 6 primary domains and 100 subdomains, with potential controls over style, format, and length, paired with inputs ranging from tens to thousands of words.
+Table 1: Comparison of existing writing benchmarks.
 Benchmark Num Domains Requirement Input Token Free Query-Form Diverse Material-Source
@@
 HelloBench 647 5 38 1,210 7,766
-WritingBench (Ours) 1,239 6 100 1,546 19,361
-Table 1: Comparison of existing writing benchmarks.
+WritingBench 1,000 6 100 1,699 19,361
 2.2 Evaluation Methods
 Using LLMs as evaluators has become a prevalent approach for evaluating the quality of generated responses.
-Typically, researchers pre-define a fixed set of evaluation dimensions applicable across all test instances Gómez-Rodríguez and Williams (2023); Marco et al.
-(2024).
-For example, SuperCLUE Xu et al.
-(2023) employs three dimensions, whereas LongBench-Write Bai et al.
-(2024) adopts six dimensions (e.g., relevance and accuracy).
-HelloBench Que et al.
-(2024) introduces task-specific dimensions, but the dimensions remain consistent across all queries of a given task.
-Although the LLM-as-a-judge approach enhances scalability, static evaluation dimensions often fail to accommodate the diversity of writing styles and specifications.
-To address this limitation, recent work Liang et al.
-(2024) trains a model to dynamically generate evaluation dimensions for individual queries.
+Typically, researchers pre-define a fixed set of evaluation dimensions applicable across all test instances (confederacy-human; pros-human-5).
+For example, SuperCLUE (superclue) employs three dimensions, whereas LongBench-Write (longwriter) adopts six dimensions.
+HelloBench (hellobench) introduces task-specific dimensions, but the dimensions remain consistent across all queries of a given task.
+Although the LLM-as-a-judge approach enhances scalability, static evaluation dimensions often fail to accommodate the diversity of writing styles and specifications, especially when dealing with inputs with enriched materials.
+To address this limitation, recent work (fennec) trains a model to dynamically generate evaluation dimensions for individual queries.
 However, the dimensions remains confined to a small predefined set.
@@
 Although existing LLMs demonstrate exceptional writing capabilities, researchers strive to further enhance their overall writing proficiency.
-Recent models, such as Weaver Wang et al.
-(2024b) and Suri Pham et al.
-(2024), have exhibited notable domain-specific strengths.
-For instance, Weaver benefits from over 200B parameter pretraining, supporting four distinct writing domains, while LongWriter Bai et al.
-(2024) specializes in length constraints.
+Recent models, such as Weaver (weaver) and Suri (suri), have exhibited notable domain-specific strengths.
+For instance, Weaver benefits from over 200B parameter pretraining, supporting four distinct writing domains, while LongWriter (longwriter) specializes in length constraints.
 However, these models experience substantial performance degradation in cross-domain scenarios and multi-constraint tasks.
@@
 3 WritingBench
-Figure 3: Domain categories in WritingBench. Inner represents the 6 primary domains and outer depicts 100 secondary subdomains (Sub = subdomains per category, Num = dataset entries).
+In this section, we will mainly introduce the four-stage human-AI collaborative construction process of WritingBench, and the query-dependent evaluation framework with its accompanied critic model for criteria-aware evaluation.
+Additionally, we train a writing-enhanced model to demonstrate the effectiveness of the framework’s data curation capability.
+3.1 Benchmark Construction
+WritingBench is constructed following a delicate pipeline (see Figure 2) that combines model-generated query refinement and human annotation, ensuring diversity and real-world relevance.
+The construction process is illustrated below.
+3.1.1 Model-Augmented Query Generation
+This phase leverages LLMs to generate an initial set of writing queries, which are then enriched and diversified through systematic guidance, with material suggestions provided as needed.
+Figure 3: Domain categories in WritingBench. Inner represents the 6 primary domains and outer depicts 100 secondary subdomains (Sub = subdomains per category).
+Table 2: Data statistics for WritingBench.
 Category Num
@@
 Domain
-Academic & Engineering 187 1,915 15,534
-Finance & Business 238 1,762 19,361
-Politics & Law 226 2,274 18,317
-Literature & Arts 242 1,133 9,973
-Education 151 1,173 10,737
-Advertising & Marketing 195 886 6,504
+Academic & Engineering 167 1,944 15,534
+Finance & Business 210 1,867 19,361
+Politics & Law 201 2,363 18,317
+Literature & Arts 183 1,266 7,675
+Education 111 1,345 10,737
+Advertising & Marketing 128 984 6,504
 Requirement
-Style 395 1,404 18,197
-Format 342 1,591 18,197
-Length 214 1,226 14,097
+Style 442 1,728 19,361
+Format 498 1,715 15,534
+Length 222 1,362 14,097
 Length
-<1K 727 443 994
-1K-3K 341 1,808 2,991
-3K-5K 94 3,804 4,966
-5K+ 77 8,042 19,361
-Table 2: Data statistics for WritingBench categorized by domain, requirement, and length.
-In this section, we will mainly introduce the construction process of our WritingBench, and the query-dependent evaluation framework with the accompanied critic model for criteria-aware evaluation.
-Additionally, we train a writing-enhanced model to demonstrate the effectiveness of our framework’s data curation capability.
-3.1 Benchmark Construction
-WritingBench is constructed following a delicate pipeline (see Figure 2) that combines model-generated query refinement and human annotation, ensuring diversity and real-world relevance.
-The two-phrase construction process is illustrated below.
-3.1.1 Model-Augmented Query Generation
-This phase leverages LLMs to generate an initial set of writing queries, which are enriched and diversified through systematic guidance, with material suggestions provided as needed.
-Phase 1: Initial Query Generation
-We begin by constructing a two-tiered domain pool grounded in real-world writing scenarios, consisting of 6 primary domains and 100 secondary subdomains (see Figure 3 and Appendix A for detailed descriptions).
-The selected domains are designed to capture both traditional and emerging user needs for AI-assisted writing, encompassing categories such as Academic & Engineering, Finance & Business, Politics & Law, Literature & Art, Education, Advertising & Marketing.
-Leveraging the domain and subdomain tags, we prompt two different LLMs (ChatGPT-4o-latest and Claude-3.5-Sonnet) to generate initial writing queries that simulate realistic user requests to maximize diversity (see Appendix B.1).
-Phase 2: Query Diversification
-To improve the diversity and practical applicability of queries, while taking into account real-world needs, we design a set of query diversification strategies (see Appendix B.2) inspired by Xu et al.
-(2024), include three core requirements and three auxiliary dimensions:
-$\bullet$ Stylistic adjustments (e.g., “Use a friendly and simple tone that kids can easily understand.”)
-$\bullet$ Format specifications (e.g., “Follow the IEEE conference template”)
-$\bullet$ Length constraints (e.g., “Generate a 500-word executive summary”)
-$\bullet$ Personalization (e.g., “From the viewpoint of an educator with two decades of experience.”)
-$\bullet$ Content specificity (e.g., “Detail the 2023 Q3 financial metrics”)
-$\bullet$ Expression (e.g., Modify the query expression to be shorter.)
-The first three categories represent typical writing capabilities, which we will evaluate through specialized assessments.
-While refining the queries, the LLM simultaneously provides material requirements (e.g., requesting financial reports as input for market analysis queries) (see Appendix B.3).
-This approach results in enriched queries accompanied by corresponding recommended reference materials.
+0 - 1K 550 470 994
+1K - 3K 292 1,832 2,991
+3K - 5K 87 3,828 4,966
+Over 5K 71 8,053 19,361
+Phase 1 - Initial Query Generation: We begin by constructing a two-tiered domain pool grounded in real-world writing scenarios, consisting of 6 primary domains and 100 subdomains (see Figure 3 and Appendix B for detailed domain statistics and descriptions; the construction process of domain system is described in Appendix A.1).
+These domains are designed to capture both traditional and emerging user needs for AI-assisted writing, categorized by topic and functionality into 6 primary domains: Academic & Engineering, Finance & Business, Politics & Law, Literature & Art, Education, and Advertising & Marketing.
+Leveraging the primary domain and subdomain tags, we prompt two different LLMs (GPT-4o and Claude-3.5-Sonnet claude-3.5)to produce an extensive pool of initial writing query drafts that simulate realistic user requests to maximize diversity (see Appendix C.2).
+Phase 2 - Query Diversification: To enhance the diversity and practical applicability of queries while addressing real-world needs, we design a set of query diversification strategies (see Appendix C.3), inspired by wizardlm.
+These strategies are divided into three core requirements and three auxiliary dimensions.
+The core requirements focus on fundamental aspects of writing: (1) Stylistic adjustments (e.g., "Use a friendly and simple tone that kids can easily understand"), (2) Format specifications (e.g., "Follow the IEEE conference template"), and (3) Length constraints (e.g., "Generate a 500-word executive summary"), which will be evaluated through specialized assessments.
+The auxiliary dimensions address additional considerations: (4) Personalization (e.g., "From the viewpoint of an educator with two decades of experience"), (5) Content specificity (e.g., "Detail the 2023 Q3 financial metrics"), and (6) Expression (e.g., "Modify the query expression to be shorter").
+While refining the queries, the LLM simultaneously provides necessary material suggestions (e.g., requesting financial reports as input for market analysis queries) (see Appendix C.4).
+This structured approach ensures that queries are both diverse and practical.
 Figure 4: Example of dynamically generating criteria for a writing query in WritingBench. Different background colors represent various types of requirements.
 3.1.2 Human-in-the-Loop Refinement
-This phase incorporates human expertise to verify model-generated queries and supplement material requirements, thereby ensuring their alignment with real-world applications and avoiding harmful content.
-Phase 1: Material Collection
-At this stage, we engage 30 trained annotators, each compensated at \$18 per hour and possessing specialized expertise proven through domain-knowledge tests.
-Their role is to collect necessary open-source materials (e.g., public financial statements or legal templates), guided by material requirements generated by LLMs.
-To minimize errors arising from parsing documents in diverse formats, the annotators carefully extract and verify the most pertinent text segments.
-Phase 2: Expert Screening & Optimization
-Next, five experts with LLM experience or industry expertise perform data screening.
-The experts conduct a delicate two-stage filtering process: (1) query adaptation: ambiguous or unrealistic queries are revised to better align with the provided materials and practical scenarios (e.g., adjusting a legal opinion query to reference specific clauses from the supplied statutes), (2) material pruning: redundant or irrelevant content is eliminated from the collected materials, ensuring that the context provided for writing tasks remained focused and relevant while guaranteeing the information is free from harmful content.
-Finally, we construct WritingBench, a benchmark comprising 1,239 queries categorized using a two-tiered taxonomy.
+The above outputs serve solely to provide a large, linguistically diverse pool of drafts.
+Human experts then verify these queries and supplement the required materials, ensuring alignment with real-world applications and avoiding harmful content.
+Phase 1 - Material Collection: At this stage, we engage 30 trained annotators, compensated at \$18 per hour and possessing specialized expertise proven through domain-knowledge tests.
+Their role is to collect open-source reference materials needed for the writing tasks (e.g., public financial statements or legal templates).
+Annotators may either adopt the sources suggested by the LLM-generated material requirements or, when more appropriate, independently identify alternative documents.
+Each query can be paired with multiple material sources, whereas in some cases no external material is required.
+To minimize errors caused by parsing documents in varied formats and to eliminate noise from irrelevant content such as web advertisements, the annotators carefully extract and verify only the most pertinent text segments.
+Phase 2 - Expert Optimization: Next, five experts with LLM experience or relevant industry expertise conduct data screening.
+The experts first exclude low-quality data that are unrealistic, overly generic, or potentially harmful.
+For the remaining entries, a rigorous two-stage review is employed: (1) query adaptation: ambiguous expressions are revised to better align with the provided materials and practical scenarios (e.g., adjusting a legal opinion query to reference specific clauses from the supplied statutes), (2) material pruning: redundant or irrelevant information is eliminated from the accompanying materials, ensuring that the context for writing tasks remains focused, relevant, and devoid of harmful content.
+Furthermore, experts are encouraged to create original, high-quality queries that are cross-validated by their peers, to further enhance diversity and realism.
+Finally, we construct WritingBench, a benchmark comprising 1,000 queries categorized using a two-tiered taxonomy, including 445 queries in Chinese and 555 queries in English.
 In comparison to existing writing benchmarks summarized in Table 1, WritingBench exhibits notable advantages in terms of the number of instances, domain diversity, requirement coverage, and variability in input lengths.
-The detailed statistical distribution of WritingBench is shown in Table 2.
+The detailed statistical distribution of WritingBench is shown in Table 3.1.1.
 3.2 Query-Dependent Evaluation Framework
-Traditional LLM-as-a-judge evaluations typically rely on fixed evaluation criteria derived from general writing assessment conventions Shao et al.
-(2024); Bai et al.
-(2024); Que et al.
-(2024).
+Traditional LLM-as-a-judge evaluations typically rely on fixed evaluation criteria derived from general writing assessment conventions (storm; longwriter; hellobench).
 However, such static criteria exhibit three critical limitations: (1) Domain exhaustiveness: fixed criteria are inadequate in adapting to specialized domains, failing to address unique characteristics found in areas like technical documentation or creative writing; (2) Requirement specificity: such criteria lack the flexibility to encompass specific requirements related to style, format, or length control; and (3) Material dependency: they are insufficient to verify whether responses appropriately utilize the provided reference materials, such as incorporating data points or maintaining narrative continuity.
-To address these challenges, we propose a query-dependent evaluation framework that enables dynamic adaptation to diverse writing scenarios. Our approach comprises two phases:
-Phase 1: Dynamic Criteria Generation
-As illustrated in Figure 4, given a query $q$ in the WritingBench, the LLM is prompted to automatically generate a set of five evaluation criteria, $C_{q}=\{c_{1},\ldots,c_{5}\}$ (opting for five criteria reflects a common approach observed in many contemporary evaluation contexts Shao et al.
-(2024); Ma et al.
-(2024); Marco et al.
-(2024)), using a carefully designed instruction to ensure structural guidance during criteria specification (see Appendix B.4).
+To address these challenges, we propose a query-dependent evaluation framework that enables dynamic adaptation to diverse writing scenarios.
+This approach comprises two phases:
+Phase 1: Dynamic Criteria Generation: As illustrated in Figure 4, given a query $q$ in the WritingBench, the LLM is prompted to automatically generate a set of five evaluation criteria, $C_{q}=\{c_{1},\ldots,c_{5}\}$ (opting for five criteria reflects a common approach observed in many contemporary evaluation contexts (storm; mops; pros-human-5)), using a carefully designed instruction to ensure structural guidance during criteria specification (see Appendix C.5).
+We utilize Claude-3.7 for generation, as it demonstrates superior diversity and comprehensiveness in criteria generation compared to models such as GPT-4o(see Appendix A.5.2 for details).
 Each criterion comprises three components: a concise name summarizing the criterion, an extended description elaborating on the evaluation focus, and detailed scoring rubrics, which provide fine-grained quality levels for the respective evaluation dimensions.
-Phase 2: Rubric-based Scoring
-For each criterion $c_{i}\in C_{q}$, the evaluator independently assigns a score on a 10-point scale to a response $r$, providing both a score and a justification.
-The final overall score is computed by averaging scores across all criteria.
-Detailed prompt is provided in Appendix B.5.
-Models Avg Languages Domains Requirements
+The generated criteria are further reviewed by human annotators to confirm their reasonableness and ensure no harmful content.
+Table 3: WritingBench performance of LLMs across 6 domains and 3 core requirements evaluated with our critic model (scale: 1-10).
+The standard deviation is computed over 3 samples.
+Domains include: (D1) Academic & Engineering, (D2) Finance & Business, (D3) Politics & Law, (D4) Literature & Art, (D5) Education, and (D6) Advertising & Marketing.
+The writing requirements assessed are: (R1) Style, (R2) Format, and (R3) Length.
+Here, “C" indicates category-specific scores.
+The latest results are available on the online leaderboard.
+Models Overall Languages Domains Requirements
 ZH EN D1 D2 D3 D4 D5 D6 R1 C R2 C R3 C
 Proprietary LLMs
-ChatGPT-4o-latest 8.16 8.3 8.1 8.1 8.1 8.2 8.1 8.4 8.1 8.3 8.7 8.2 8.9 8.2 8.3
-o1-Preview 8.15 8.1 8.2 8.0 8.1 8.2 8.2 8.4 8.1 8.2 8.6 8.2 8.8 8.2 8.2
-Claude-3.5-Sonnet 7.71 7.7 7.7 7.6 7.5 7.6 7.7 7.9 8.0 7.9 8.5 7.7 8.5 7.9 8.0
-Gemini-1.5-Pro 7.78 7.8 7.7 7.7 7.5 7.8 7.9 8.0 7.9 7.9 8.6 7.9 8.8 7.9 8.0
-Qwen-Max 8.37 8.4 8.3 8.3 8.3 8.4 8.4 8.5 8.4 8.5 8.7 8.4 9.0 8.4 8.5
+Claude-3.7-thinking 7.91±0.111 7.9 7.9 7.9 7.8 7.8 8.0 8.0 8.1 7.9 8.7 8.0 8.4 8.0 8.1
+Claude-3.7 7.85±0.110 7.9 7.8 7.8 7.8 7.7 7.9 8.0 8.1 7.9 8.6 7.9 8.3 8.0 8.1
+Qwen-Max 7.16±0.041 7.2 7.1 7.1 6.9 7.0 7.3 7.4 7.5 7.2 8.3 7.3 7.8 7.2 7.5
+o1-Preview 6.89±0.039 6.8 7.0 6.9 6.8 6.7 7.0 7.1 7.2 6.9 8.0 7.0 7.5 7.1 7.3
+GPT-4o 6.81±0.028 6.9 6.7 6.8 6.6 6.7 6.8 7.0 7.1 6.9 8.0 7.0 7.5 6.8 6.8
+Gemini-1.5-Pro 6.21±0.018 6.2 6.2 6.2 5.8 6.0 6.4 6.6 6.7 6.2 7.2 6.4 7.1 6.4 6.0
 Open-source LLMs
-Deepseek-R1 8.55 8.7 8.5 8.5 8.5 8.6 8.6 8.7 8.6 8.7 8.9 8.6 9.0 8.6 8.7
-Deepseek-V3 7.95 8.0 7.9 7.9 7.8 8.0 7.8 8.2 8.0 8.1 8.6 8.0 8.9 8.0 8.2
-Mistral-Large-Instruct 7.64 7.6 7.7 7.7 7.6 7.8 7.3 7.9 7.6 7.7 8.2 7.7 8.7 7.7 7.9
-Qwen-2.5-72B-Instruct 7.90 8.0 7.9 8.0 7.8 8.1 7.7 8.2 7.8 8.0 8.3 8.0 8.8 7.9 8.0
-Qwen-2.5-7B-Instruct 7.43 7.3 7.5 7.7 7.4 7.6 6.9 7.8 7.3 7.5 7.9 7.6 8.6 7.4 7.5
-Llama-3.3-70B-Instruct 7.01 6.7 7.3 7.0 6.9 7.0 6.8 7.3 7.3 7.1 7.8 7.1 8.2 7.0 7.2
-Llama-3.1-8B-Instruct 6.35 5.7 6.9 6.6 6.4 6.1 6.0 6.7 6.6 6.4 7.0 6.4 7.6 6.3 6.4
+Deepseek-R1 7.70±0.053 8.0 7.5 7.6 7.4 7.6 7.8 7.8 8.1 7.7 8.4 7.9 8.3 7.7 7.5
+Deepseek-V3 6.35±0.022 6.3 6.4 6.4 6.1 6.2 6.3 6.6 6.8 6.4 7.6 6.5 7.1 6.5 6.4
+Mistral-Large-Instruct 6.00±0.076 5.9 6.1 6.2 5.9 5.9 5.7 6.4 6.4 6.1 7.3 6.1 6.5 6.0 6.0
+Qwen-2.5-72B-Instruct 6.40±0.061 6.4 6.4 6.6 6.2 6.4 6.2 6.7 6.6 6.5 7.7 6.5 6.9 6.5 6.5
+Qwen-2.5-7B-Instruct 5.64±0.083 5.5 5.8 5.9 5.6 5.6 5.1 6.1 5.9 5.7 7.0 5.7 6.1 5.6 5.6
+Llama-3.3-70B-Instruct 5.05±0.011 4.5 5.5 5.1 4.9 4.8 4.8 5.3 5.6 5.0 5.0 5.1 5.9 5.1 5.0
+Llama-3.1-8B-Instruct 4.42±0.004 3.7 5.0 4.1 4.4 4.0 4.1 4.7 5.0 4.4 4.4 4.5 5.3 4.4 4.3
 Capability-enhanced LLMs
-Suri 4.97 4.4 5.5 5.6 5.3 5.0 4.1 5.0 5.1 4.8 5.2 5.0 5.4 4.5 4.0
-LongWriter 7.91 7.9 7.9 8.0 8.1 8.1 7.7 8.1 7.6 7.9 8.2 8.1 8.8 7.7 7.7
-Qwen-2.5-7B-filtered 8.49 8.6 8.4 8.4 8.4 8.6 8.4 8.6 8.5 8.6 8.8 8.5 9.0 8.5 8.6
-Llama-3.1-8B-filtered 8.49 8.6 8.4 8.5 8.4 8.6 8.4 8.6 8.5 8.6 8.8 8.5 8.9 8.5 8.5
-Table 3: WritingBench performance of different LLMs across 6 domains and 3 writing requirements evaluated with our critic model (scale: 1-10).
-The six domains include: (D1) Academic & Engineering, (D2) Finance & Business, (D3) Politics & Law, (D4) Literature & Art, (D5) Education, and (D6) Advertising & Marketing.
-The three writing requirements assessed are: (R1) Style, (R2) Format, and (R3) Length.
-Here, “C" indicates category-specific scores.
-3.3 Critic Model
+Suri 3.20±0.042 2.5 3.8 3.6 3.5 3.0 2.5 3.2 3.6 3.2 3.7 3.1 3.2 3.0 3.0
+LongWriter 6.27±0.081 6.2 6.4 6.4 6.4 6.3 6.0 6.5 6.0 6.3 7.4 6.3 6.7 6.3 6.8
+Qwen-2.5-7B-filtered 7.44±0.058 7.7 7.2 7.4 7.2 7.5 7.3 7.7 7.7 7.5 8.4 7.6 8.1 7.4 7.2
+Llama-3.1-8B-filtered 7.39±0.045 7.5 7.3 7.4 7.2 7.3 7.3 7.5 7.8 7.4 8.3 7.5 8.0 7.4 7.1
+We identify three commonly encountered writing requirements: style, format, and length from real-world writing scenarios (frequency analysis detailed in Appendix A.1).
+For each requirement, we create two evaluation subsets.
+Two annotators review each query to classify each criteria in $C_{q}$ into style, format, length, or none.
+Based on their annotations, we perform final verification and define the subsets as follows: (1) The first requirement-involved subset includes all queries involving a specific requirement and their entire set of criteria.
+For example, if any criterion in $C_{q}$ relates to style, the query and its entire criteria set $C_{q}$ belong to this subset.
+Scores for all five criteria are calculated, corresponding to the R1/R2/R3 columns in Table 3.
+(2) The second category-specific subset includes only the criteria related to the specific requirement.
+For instance, if $c_{1}$ and $c_{4}$ in $C_{q}$ are format-related, then the query and the subset $c_{1},c_{4}$} belong to this subset.
+Only the scores for $c_{1},c_{4}$} are calculated, corresponding to the C columns.
+The first subset serves as an overall evaluation of writing quality for a specific requirement, while the second subset provides more targeted insights into performance on that particular capability.
+Phase 2 - Rubric-based Scoring: For each criterion $c_{i}\in C_{q}$, the evaluator independently assigns a score on a 10-point scale to a response $r$, along with a justification.
+The final overall score is computed by averaging scores across all criteria.
+Scoring prompt is provided in Appendix C.6.
 To alleviate the computational overhead with LLM-based evaluation, we develop a dedicated critic model, $\mathcal{M}$, designed to implement our rubric-based scoring framework.
 Specifically, this model performs the mapping $\mathcal{M}_{c}:(q,r,c_{i})\mapsto[1,10]\times\mathcal{J}$, where the output consists of a numerical score and corresponding justification text, $\mathcal{J}$, both in accordance with the predefined evaluation rubric.
-We fine-tune the critic model on a dataset comprising 50K instances, which are collected using LLMs in our experiments.
-The dataset encompasses diverse queries, evaluation criteria, and model responses to enhance the robustness of evaluation.
-The experiments presented in Section 4.3 validate the consistency of the critic model.
+The critic model is fine-tuned on a dataset of 155K instances scored by Claude-3.7, which demonstrates higher human alignment consistency (as discussed in Section 4.3).
+When building the training data of scoring prompts, the queries and criteria are drawn from WritingBench.
+The response portion is generated using approximately 40 different models, including those evaluated in Section 4.2, as well as additional models of varying types and sizes, such as Claude-3.5-Haiku claude-3.5 and Llama-3.2-1B-Instruct llama-3.2.
+Finally, after removing instances where scoring failed, approximately 155K samples are constructed for training.
+The experiments described in Section 4.3 confirm the consistency of the critic model.
 3.3 Evaluation-Guided Data Curation for Writing Enhancement
 The query-dependent evaluation framework enables systematic data curation across diverse writing tasks through two-phase filtering.
-First, applying instance-specific criteria generation to initial 24K SFT samples.
-Subsequently, the critic model filters 50% samples, yielding 12K framework-curated samples with higher quality.
+To validate its effectiveness in data curation, we conduct SFT experiments using constructed datasets.
+We utilize the secondary domain taxonomy of WritingBench and follow the first two steps outlined in Section 3.1.1 (removing the material generation suggestion part from the prompt in Appendix C.4, requiring only the refined query from the model) to let LLMs generate writing queries.
+To enrich query diversity, we employ two models, GPT-4o and Claude-3.5-Sonnet, to each generate 60 queries in both Chinese and English for every subdomain, totaling 24K queries.
+Responses are uniformly generated using an advanced model, Deepseek-R1.
+Subsequently, we apply the query-dependent evaluation metric with our critic model described in Section 3.2 to select top-50% samples per subdomain, resulting in 12K high-quality data samples.
 Fine-tuning experiments are conducted using the Llama-3.1-8B-Instruct and Qwen-2.5-7B-Instruct models.
-Evaluations are performed on two writing benchmarks to validate the framework’s effectiveness in identifying high-quality writing samples.
+Both models (Qwen-2.5-7B-filtered and Llama-3.1-8B-filtered) demonstrated significant performance improvements over their base versions and even outperformed larger models such as Llama-3.3-70B-Instruct and Qwen-2.5-72B-Instruct in our experiments on two benchmarks, confirming the framework’s effectiveness in identifying high-quality writing samples.
 4 Experiment
-Figure 5: Scores of different models across different subdomains in WritingBench.
-Red indicates higher score and blue refers to lower score.
-The figures are the average score of each subdomain for different models.
+Figure 5: Score heatmap of models across 100 subdomains in WritingBench. The figure shows the average score per subdomain for each model. Red indicates higher score and blue indicates lower.
 4.1 Experiment Settings
-This section outlines the settings employed in our experiments using the WritingBench framework. The evaluation protocol, and the model training configurations are as follows:
-Evaluation Protocol: Uniform generation settings are applied to model responses on WritingBench, allowing up to 16,000 tokens.
-We set the temperature at 0.7, with a top-k of 20 and top-p of 0.8, ensuring a balance of creativity.
-The critic model uses the same parameters, with input length capped at 2,048 tokens for scoring stability.
+This section outlines the settings employed in our experiments using the WritingBench framework. The evaluation protocol and the model training configurations are as follows:
+Evaluation Protocol: Uniform generation settings are applied to model responses on WritingBench, with maximum sequence lengths capped at 16,000 tokens (or platform maximum if lower).
+We configure the generation process with a temperature of 0.7, a top-k value of 20, and top-p of 0.8, ensuring a balance of creativity and coherence.
+For both the critic model and LLM scoring, we use consistent configurations.
+The temperature is set to 1.0, in line with the Arena-Hard-Auto framework arena-hard.
+We apply a top-p value of 0.95 without any top-k filtering, and the maximum length is set at 2,048 tokens to ensure efficient and reliable output generation.
 Model Training: The critic model is fine-tuned on the Qwen-2.5-7B-Instruct base model, using the AdamW optimizer with a 7e-6 learning rate.
-It is trained on 50K SFT data for 3 epochs across 8xA100 GPUs (batch size 64, 8-step accumulation).
+It is trained on 155K SFT data for 3 epochs across 16xA100 GPUs (batch size 64, 4-step accumulation).
 The writing models, trained on both Qwen-2.5-7B-Instruct and Llama-3.1-8B-Instruct for 5 epochs on 32xA100 GPUs, achieving a total batch size of 128 with 4-step gradient accumulation.
+Table 4: Comparison of human consistency scores across different criteria generation methods. Claude corresponds to Claude-3.7. Critic refers to the critic model.
+Evaluation Metric Judge Score
+Static Global GPT-4o 69%
+Static Domain-Specific GPT-4o 40%
+Dynamic Query-Dependent GPT-4o 79%
+Static Global Claude 67%
+Static Domain-Specific Claude 58%
+Dynamic Query-Dependent Claude 87%
+Dynamic Query-Dependent Critic 84%
+Table 5: Performance of our writing models on two benchmarks. ‘-filtered’ indicates models trained with filtered data, while ‘-all’ uses the full dataset. Eval2 refers to LongBench-Write longwriter.
+Models WritingBench Eval2
+Deepseek-R1 7.70 4.79
+Qwen-2.5-7B-Instruct 5.64 4.39
+Llama-3.1-8B-Instruct 4.42 3.12
+Qwen-2.5-7B-all 7.36 4.69
+Qwen-2.5-7B-filtered 7.44 4.70
+Llama-3.1-8B-all 7.34 4.65
+Llama-3.1-8B-filtered 7.39 4.65
 4.2 Comparison between LLMs
-We evaluate 16 LLMs: ChatGPT-4o-latest Hurst et al.
-(2024), o1-Preview Jaech et al.
-(2024), Claude-3.5-Sonnet Anthropic (2024), Gemini-1.5-Pro Reid et al.
-(2024), Qwen-Max Team (2025), Deepseek-R1 DeepSeek-AI et al.
-(2025), Deepseek-V3 DeepSeek-AI et al.
-(2024), Mistral-Large-Instruct MistralAI (2024), Qwen-2.5-72B-Instruct and Qwen-2.5-7B-Instruct Yang et al.
-(2024), Llama-3.3-70B-Instruct and Llama-3.1-8B-Instruct Dubey et al.
-(2024), Suri Pham et al.
-(2024), LongWriter Bai et al.
-(2024), and our writing model, Qwen-2.5-7B fine-tuned model and Llama-3.1-8B fine-tuned model, on WritingBench.
-Each query is assessed by 5 instance-specific criteria using a 10-point scale, with subdomain-specific subcategory heatmaps revealing detailed variations.
-Key Insights from Domain Scores: From Table 3, we observe that Education (D5) always yields high performance among models, followed by Academic & Engineering (D1).
-Literature & Art (D5) is the lowest-scoring domain, showing notable variance.
-CoT-capable models like Deepseek-R1 and o1-preview outperform their non-CoT counterparts models, indicating CoT’s potential in handling narrative and creative content.
-Figure 5 identifies challenges in niche areas such as novel continuations, tender proposals, and white papers where models generally score lower.
+We evaluate 17 LLMs: GPT-4o [^1] In this paper, we specifically use GPT-4o, version gpt-4o-2025-01-29.
+(4ol), o1-Preview (o1), Claude-3.7 and its thinking version Claude-3.7-thinking (claude), Gemini-1.5-Pro (gemini), Qwen-Max (qwen-max), Deepseek-R1 (r1), Deepseek-V3 (deepseek-v3), Mistral-Large-Instruct (mistral), Qwen-2.5-72B-Instruct and Qwen-2.5-7B-Instruct (qwen2.5), Llama-3.3-70B-Instruct and Llama-3.1-8B-Instruct (llama), Suri (suri), LongWriter (longwriter), and our writing model, Qwen-2.5-7B-filtered (fine-tuned on Qwen-2.5-7B-Instruct) and Llama-3.1-8B-filtered (fine-tuned on Llama-3.1-8B-Instruct), on WritingBench.
+For models capable of reasoning, the reasoning content is excluded from the evaluated response.
+Each query is assessed by 5 instance-specific criteria using a 10-point scale scoring by our critic model.
+The overall scores of the models are presented in Table 3, along with their performance across two languages and various domain and requirement subsets.
+Detailed variations are further revealed through subdomain-specific subcategory heatmap in Figure 5.
+Key Insights from Domain Scores: The Education (D5) and Advertising & Marketing (D6) domains always yield substantial performance across models.
+In contrast, the Academic & Engineering (D1) and Finance & Business (D2) domains present greater challenges, as these tasks inherently require more sophisticated information processing and integration capabilities.
+Our evaluation across 100 subdomains further identifies persistent difficulties in niche areas like writing bid proposals, financial reports, and white papers, where models generally achieve lower scores.
 These tasks demand a higher level of knowledge, long-text generation capabilities, and adherence to contextual consistency, pinpointing areas for further enhancement.
-Key Insights from Requirement Scores: Table 3 presents scores for three requirement subsets: R columns showing overall scores and C columns showing requirement-specific averages.
-For example, if a style query includes two relevant criteria (as confirmed by human annotation), the C column averages scores for those criteria only.
-Style (R1) and Format (R2) dimensions reveal model distinctions , with Deepseek-R1 and Qwen-Max often excels others.
-Nevertheless, length requirements (C of R3) remains challenging, especially in section-specific constraints (see Appendix 16 for examples) and extended text production (see Section 4.5).
-The overall analysis of WritingBench experiment highlights: Deepseek-R1 consistently leads across both domain and requirement dimensions, showcasing its versatility and strong language capabilities.
-There is significant performance variance within creative content domains, with CoT models showing notable promise—a hypothesis further confirmed by our comparison (see Appendix C.1, where the model incorporating CoT mechanisms outperform the one without CoT on creative writing tasks).
-Additionally, we observe that advanced models often achieve higher specialized scores (C column) for the three common requirements than their overall scores (R column).
-The criteria outside these specialized sets tend to emphasize content aspects, such as integration with source materials and writing depth, highlighting the need to improve content quality.
-This analysis benchmarks existing model capabilities and underscores specific areas needing improvement for future development.
+The Literature & Art (D4) domain exhibits notable performance variance among models.
+Reasoning-capable architectures such as Claude-3.7-thinking, Deepseek-R1, and o1-Preview outperform their non-reasoning counterparts, indicating the potential of CoT techniques in processing narrative and creative content.
+To further validate CoT’s efficacy in creative writing, we employ the 12k SFT dataset described in Section 3.3 to fine-tune Qwen-32b-Instruct using both CoT-integrated and non-CoT approaches.
+Evaluations are conducted on the Literature & Art (D4) subset of WritingBench and EQ-Bench eqbench, a specialized benchmark designed for creative writing tasks.
+The results reveal that the reasoning model consistently surpass both the baseline Qwen-32b-Instruct and its non-reasoning variants, with detailed results provided in Appendix A.2.
+Key Insights from Requirement Scores: Most models perform well in the style dimension, followed by format, with length being the weakest.
+We observe that advanced models often achieve higher specialized scores (C column) for the three common requirements compared to their overall scores (R column).
+Criteria outside these specialized sets tend to emphasize content-related aspects, such as integration with source materials and writing depth, underscoring the need to improve content quality.
+Length requirements remain particularly difficult, especially in section-specific constraints (see Appendix 7 for examples) and extended text generation.
+Additionally, we evaluate the performance of LLMs across varying input and output lengths (see Appendix A.4 for details).
+Advanced models generally sustain consistent performance across varying input lengths, leveraging their strong long-context comprehension abilities.
+However, regarding output length, most models demonstrate inherent limitations, typically capping their responses at around 3,000 tokens.
+In contrast, Claude-3.7 and its reasoning-enhanced version, as well as Qwen-Max, stand out for their capacity to generate extended outputs effectively.
+These results emphasize the critical need to improve long-output generation and refine length optimization in writing-related tasks.
+The overall analysis of WritingBench experiment highlights: (1) Claude-3.7-thinking consistently leads across both domain and requirement dimensions, followed by its non-reasoning variant, showcasing versatility and strong language capabilities; (2) A significant performance variance is observed within creative content domains, where models incorporating CoT mechanisms surpass those without, demonstrating CoT’s potential in LLM writing; (3) Cross-lingual inconsistencies in models like Deepseek-R1 and Llama-3.3-70B-Instruct suggest limitations in multilingual knowledge alignment.
+Furthermore, detailed analysis on subdomains reveals persistent difficulties in knowledge-intensive tasks.
+Ablation experiments on output length emphasize that generating long-form outputs remains an obstacle for current models.
+This analysis not only benchmarks the existing capabilities of these models but also underscores specific areas needing improvement for future development.
+For the most up-to-date results, please refer to the online leaderboard [^2] For clearer display, online leaderboard scores use the same calculations but are rescaled to a 100-point scale..
 4.3 Human Consistency
-To validate the alignment between automated evaluation and human judgment, we conducted human evaluation on 300 queries, covering all 100 subdomains.
+To validate the alignment between automated evaluation and human judgment, we conduct a human evaluation study involving 300 queries (constructed in the same pipeline as described in Section 3.1 but not included in WritingBench).
 Five professionally trained annotators with linguistic backgrounds perform pairwise comparisons of model responses.
-For each query, two responses are randomly selected from different models.
-Annotators select the preferred response or declared equivalence based on the query’s requirements.
-We compare our dynamic query-dependent criteria against two baselines: static globally uniform criteria and static domain-specific customized criteria (static criteria are designed by domain experts) on two LLM judges: ChatGPT-4o-latest and Claude-3.5-Sonnet.
-Evaluation Metric Judge Score
-Static Global ChatGPT 69%
-Static Domain-Specific ChatGPT 40%
-Dynamic Query-Dependent ChatGPT 79%
-Static Global Claude 65%
-Static Domain-Specific Claude 59%
-Dynamic Query-Dependent Claude 87%
-Dynamic Query-Dependent Critic Model 83%
-Table 4: Comparison of human consistency scores across different criteria generation methods. ChatGPT corresponds to ChatGPT-4o-latest, Claude corresponds to Claude-3.5-Sonnet.
-As shown in Table 4, our dynamic query-dependent criteria achieve superior human alignment compared to static, both globally uniform or domain-specific customized criteria.
+For each query, two responses are generated by two randomly selected models (drawn from the same set of 17 models used for evaluation in Section 4.2).
+The triplet <Query, Response_A, Response_B> is presented on the same page to enable direct comparison.
+Annotators are instructed to carefully read the query and select their preferred response or declare equivalence (A/B/Tie), with no criteria information provided to ensure unbiased decision-making (detailed instructions and the annotation interface are provided in Appendix A.3).
+We then compare these human preferences with the scores assigned by our critic model to the responses — if the critic model rate Response_A higher and the annotators also prefer Response_A, it is counted as alignment.
+To further evaluate the effectiveness of our dynamic query-dependent criteria, we compare it against two baselines: static globally uniform criteria and static domain-specific customized criteria (designed by domain experts).
+This comparison was conducted using two LLM-based judges, GPT-4o and Claude-3.7.
+Both models score the responses using the rubric-based scoring method described in Section 3.2, using scoring prompts provided in Appendix C.6.
+As shown in Table 4.1, our dynamic query-dependent criteria achieve superior human alignment compared to static, both globally uniform or domain-specific customized criteria.
 Notably, domain-specific criteria underperform despite customization, since our queries involve highly diverse tasks and varied sources.
 These findings confirm that context-sensitive query-dependent evaluation better captures real-world writing complexity compared to conventional static approaches.
-Furthermore, the critic model attains 83% agreement, confirming its practical viability.
+Furthermore, the critic model attains 84% agreement, confirming its practical viability.
 4.4 Ablation of Data Curation for Writing-Enhanced Models
-Leveraging the query-dependent evaluation framework described in Section 3.2, we perform instance-specific criteria generation an initial 24K SFT dataset spanning various domains, including long-input queries enriched with supplemental materials.
-Our critic model then apply a query-dependent evaluation to filter this dataset down to a high-quality 12K subset.
-We conduct fine-tuning experiments using the Llama-3.1-8B-Instruct and Qwen-2.5-7B-Instruct models.
-As shown in 5, both models, trained on the filtered data, demonstrate significant performance improvements over their previous versions.
-Notably, they outperform models trained on the full dataset and even approached the capabilities of SOTA models.
-Further validation on an alternative benchmark, LongBench-Write (following the settings of quality evaluation in Bai et al.
-(2024)), confirm these findings, with our models surpassing baseline performance.
-These results validate the robustness of our query-dependent evaluation strategy and the utility of our critic model for curating high-quality writing samples.
-This integrated approach enables smaller models to compete with, and even surpass, larger models across diverse writing tasks.
-Models WritingBench Benchmark2
-Deepseek-R1 8.55 4.79
-Qwen-2.5-7B-Instruct 7.43 4.39
-Llama-3.1-8B-Instruct 6.35 3.12
-Qwen-2.5-7B-all 8.46 4.69
-Qwen-2.5-7B-filtered 8.49 4.70
-Llama-3.1-8B-all 8.45 4.65
-Llama-3.1-8B-filtered 8.49 4.65
-Table 5: Performance evaluation of our writing models on two benchmarks.’-filtered’ indicates models trained with filtered data, while ’-all’ refers to those trained with the full dataset.
-4.5 Ablation of Length
-We assess the performance of LLMs across varying input and output lengths, with statistical validity ensured by excluding intervals containing fewer than 5 samples.As demonstrated in Appendix 6, experiments reveal that most SOTA models generally maintain consistent performance regardless of input length variations, attributable to their advanced long-context comprehension capabilities.
-However, analysis in Appendix 7 shows that most models exhibit inherent limitations in response generation length, typically producing outputs constrained to approximately 3,000 tokens.
-While output quality remains relatively stable within this range, smaller models suffer more performance degradation characterized by repetitive outputs.
-Notably, only two models, Qwen-Max and LongWriter, which incorporate architectural optimizations for long-form generation, effectively support extended response lengths.
-It underscores the importance of enhancing long-output capabilities and length optimization in writing.
+To validate the data curation capabilities of the query-dependent evaluation framework described in Section 3.2, we conduct fine-tuning experiments on two datasets: the initial 24K dataset constructed in Section 3.3 and the 12K subset curated using the query-dependent evaluation framework.
+We experiment with two models of different architectures, Llama-3.1-8B-Instruct and Qwen-2.5-7B-Instruct, and evaluate their performance on two benchmarks: WritingBench and LongBench-Write, a general-purpose writing benchmark (following the quality evaluation settings outlined in (longwriter)).
+As shown in Table 4.1, both models trained on the filtered 12K dataset demonstrate significant performance improvements over their previous versions.
+Notably, they outperform models trained on the full 24K dataset and even approach the capabilities of advanced models.
+These results validate the robustness of our query-dependent evaluation strategy and highlight the effectiveness of our critic model in curating high-quality writing samples.
+This integrated approach enables smaller models to compete with, and in some cases surpass, larger models across a wide range of writing tasks.
 5 Conclusion
 In this paper, we introduce WritingBench, a comprehensive benchmark designed to evaluate LLMs in generative writing across diverse domains.
-It includes 1,239 queries spanning 6 primary domains and 100 subdomains, providing evaluation dimensions for style, format, and length requirements.
+It includes 1,000 queries spanning 6 primary domains and 100 subdomains, providing evaluation dimensions for style, format, and length requirements.
 Our query-dependent evaluation framework, supported by a critic model, achieves high human alignment.
-Evaluation efficiency is further demonstrated by compact models trained on curated data, achieving performance comparable to SOTA models.
+Evaluation efficiency is further demonstrated by compact models trained on curated data, outperforming GPT-4o in writing.
 By making WritingBench and its resources publicly available, we aim to foster further research and advancements in LLM writing capabilities.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: The methodologies for the contributions mentioned in the abstract and introduction are detailed in the subsections of Section 3, with experimental validation provided in Section 4.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
     Limitations
-This work faces three primary limitations that warrant consideration.
-First, both the writing model and critic model are primarily trained using conventional SFT approaches, omitting systematic exploration of enhanced optimization strategies.
-While we demonstrate the partial efficacy of the CoT mechanisms in creative domains, their potential remains unexplored compared to established successes in mathematical reasoning tasks.
-Second, our evaluation framework exhibits diminished precision in handling complex multi-dimensional length requirements, including temporal sequencing constraints and section-specific word counts.
-This limitation underscores the necessity for enhanced scoring methodologies that integrate learned metrics with structured rule-based evaluations to better regulate output specifications.
-Third, inherent challenges persist in obtaining reliable pairwise preference annotations for compositional tasks.
-Despite rigorous annotation protocols, human evaluators inevitably introduce subjective biases when assessing two fair-well responses, particularly regarding narrative preferences and contextual interpretations.
-While our consensus-building procedures mitigate some variability, absolute alignment with diverse user preferences remains theoretically unattainable.
-Appendix A Benchmark Statistics
-A.1 Overview of Six Primary Domains
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: Limitations are discussed in Appendix D.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [Yes]
+  14. Justification: Section 4 and Appendix A provide detailed supporting experiments and analyses for the conclusions drawn in the paper.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19.
+Justification: Section 3.1 of this paper provides a detailed description of the benchmark construction pipline with relevant prompts listed in Appendix C.
+The evaluation dataset and evaluation code can be accessed in the GitHub Repo: https://github.com/X-PLUG/WritingBench.
+The training parameters and datasets for the critic model and writing model are illustrated in Section 4.1, Section 3.3 and Section 3.2.
+The models have also been open-sourced on Hugging Face (download links can be found in the GitHub repository).
+The latest leaderboard evaluation results can be found at Leaderboard: https://huggingface.co/spaces/WritingBench/WritingBench.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: The benchmark dataset and code have been open-sourced in the GitHub Repo: https://github.com/X-PLUG/WritingBench with detailed instructions provided in README.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: The benchamrk evaluation settings and training parameters and datasets for the critic model and writing model are provided in Section 4.1, Section 3.3 and Section 3.2.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34. Justification: We provide standard deviation for benchmark evaluation in Table 3.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: The compute resources are provided in Section 4.1.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: The research conducted in the paper conforms with the NeurIPS Code of Ethics.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. Justification: The impacts are discussed in Appendix E.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [Yes]
+  54. Justification: Our benchmark serves exclusively for evaluation purposes. All released prompts undergo manual review to eliminate harmful information.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: The assets used in the paper have been properly cited.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [Yes]
+  64. Justification: We have fully discussed the benchmark in this paper and provided instructions and terms of use in GitHub Repo: https://github.com/X-PLUG/WritingBench.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [Yes]
+  69. Justification: The instructions and compensation for human annotators are provided.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [Yes]
+  74.
+Justification: We have adhered to local laws regarding ethical approval.
+All annotators involved are employed by a professional annotation company, and they have been informed of the data usage and provided consent.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [Yes]
+  79. Justification: The involvement of LLMs in the benchmark construction process and scoring is described in Section 3.1 and Section 3.2.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
+Appendix A Experiment Results
+A.1 Validation of Domain Taxonomy Construction and Requirement Dimension
+With existing benchmarks focusing on limited tasks (e.g.
+fiction), the domain taxonomy of our benchmark is rigorously designed to reflect real-world writing scenarios.
+We leverage the industrial background of team members to refine the domain taxonomy using over 200K anonymized real-user writing query data under strict data security protocols.
+Our initial two-tier domain system is established, referencing the mature workflows of industry product teams, with an additional “Other” category under each primary domain.
+This is iteratively refined through multiple rounds: in each iteration, we use the prompt outlined in Appendix C.1 and perform classification using GPT-4o to tag 2K randomly selected writing queries data according to the current subdomain labels, with each query potentially receiving multiple tags.
+Subsequently, human annotation validate the tags and adjust the domain settings for the next iteration, aiming to reduce overlaps and minimize the “Other” category.
+The final tags achieve stable after three iterations, ensuring broad domain coverage.
+To evaluate the classification accuracy of GPT-4o, we validate the effectiveness of the model’s automatic annotations on a writing task tag system provided by the production team.
+This tag system consists of 8 primary domains and 44 subdomains, with a dataset of 3,000 entries that have been manually annotated and verified under strict data security protocols.
+The recall rate is calculated such that if a manually assigned tag is found within the list of tags generated by the model, it is considered successfully retrieved.
+The overall recall rate for primary domains is approximately 96%, while the recall rate for subdomains is about 93%.
+By analyzing the queries provided by the production team, we identify three prevalent requirement dimensions: style, format, and length.
+To validate these categories, we perform model-based annotations across several iterations, using the same prompt outlined in Appendix C.1.
+After multiple rounds of analysis, the average frequencies are determined as follows: style-related requirements account for approximately 29.4%, format-related requirements comprise about 22.48%, and length-related requirements make up around 20.8%.
+The remaining requirements largely correspond to criteria closely associated with writing materials or other demands that are challenging to generalize into abstract dimensions.
+A.2 Ablation of CoT in Creative content
+To validate the impact of CoT reasoning on creative content generation, we conduct SFT experiments on filtered subset described in Section 3.3 on Qwen-2.5-32B-Instruct.
+Two variants are developed: 1) A base model trained with CoT-formatted instructions, adhering to DeepSeek-R1’s original output format, and 2) an ablated version (-w/o CoT) trained using only the response content.
+These models are evaluated on the Literature & Art (D4) subset of WritingBench and EQBench, a specialized benchmark for creative writing evaluation, to thoroughly examine their capabilities in generating creative content.
+Table 6: Ablation of CoT in creative content on two benchmarks.
+Models WritingBench-D4 EQBench
+Deepseek-R1 7.70 84.99
+Qwen-2.5-32B-Instruct 5.59 48.17
+Qwen-2.5-32B-CoT 7.58 82.48
+- w/o CoT 7.54 79.43
+As shown in Table 6, the CoT-enhanced model outperform the non-reasoning ones on both WritingBench-D4 and EQBench, showing CoT’s effectiveness in generating creative content.
+A.3 Human Consistency
+Human evaluations are conducted via DingTalk Docs (a cloud-based collaborative platform [^3] https://www.dingtalk.com/en), where annotators access standardized spreadsheets.
+The spreadsheets contain triplets of <Query, Response_A, Response_B> with queries highlighted in yellow and instructions are provided in blue header sections (refer to the interface screenshot in Figure 6).
+To mitigate positional bias, response ordering is randomized per instance, and annotators could zoom into cells for full-text review.
+The interface enable keyword searches across responses and queries for systematic comparison.
+The test set excludes inputs exceeding 5,000 tokens to reduce cognitive load, with explicit instructions emphasizing content quality over text length or surface formatting.
+Five linguists achieved substantial agreement ($\kappa=0.69$), demonstrating rigorous annotation reliability.
+Figure 6: Annotation Interface.
+A.4 Ablation of Length
+We assess the performance of LLMs across varying input and output lengths (see Figure 7), with statistical validity ensured by excluding intervals containing fewer than 5 samples.
+Experiments on input length reveal that most SOTA models generally maintain consistent performance regardless of input length variations, attributable to their advanced long-context comprehension capabilities.
+However, analysis on output length shows that some models exhibit inherent limitations in response generation length, typically producing outputs constrained to approximately 3,000 tokens.
+Small models, such as Suri, Qwen-2.5-7B-Instruct, Llama-3.1-8B-Instruct, suffer more performance degradation characterized by repetitive outputs.
+Notably, Claude-3.7 and its reasoning model, Qwen-Max and LongWriter effectively support extended response lengths.
+These findings highlight the importance of improving long-output generation capabilities and optimizing length handling in writing tasks.
+Figure 7: Scores of different models across various input and output lengths on the WritingBench.
+A.5 Case Study
+A.5.1 Bad Case of Length Constraint
+Current models often face challenges when handling length-related requirements. Figure 7 illustrates bad cases where the LLM struggle with adhering to length constraints.
+Table 7: Bad case of length requirement. Query from WritingBench with index 539, response generated by Gemini-1.5-Pro.
+Query Create advertising slogans for a pet grooming salon, including: - Main slogan (10-15 words) - Sub-slogan (15-20 words) - Service tagline (8-12 words) - Brand promise (10-15 words) The slogans should reflect professional pet grooming and care services while emphasizing thoughtful care for pet owners.
+The slogans should be catchy, memorable, with warm and friendly wording, and a light, cheerful tone.
+Response **Main Slogan:** Pamper Your Pet, Delight Your Heart.
+*Sub-Slogan:** Expert grooming and loving care, making your furry friend feel fabulous, inside and out.
+*Service Tagline:** Gentle paws, happy pets, stress-free smiles.
+*Brand Promise:** Quality grooming, compassionate care, every time.
+A.5.2 Criteria Generation Comparison
+We compare the ability of various models to generate criteria and ultimately selected Claude-3.7, which demonstrates advantages in diversity, comprehensiveness, and rationality of the criteria.
+Figure 8 presents a comparison between Claude-3.7 and GPT-4o based on the same example query.
+The criteria descriptions generated by Claude-3.7 show a higher degree of integration with the specific requirements of the query.
+In contrast, we observe in more cases that GPT-4o tends to uses similar criteria and has a lower level of integration with the query, such as not adequately considering information about the material.
+Figure 8: Compasion of criteria generated by Claude-3.7 and GPT-4o.
+Appendix B Benchmark Statistics
+B.1 Overview of Six Primary Domains
   1. 1.
@@
 The writing is dynamic and strategic, with a creative twist, necessitating a deep understanding of audience psychology and trend dynamics to ensure that the content is appealing and strategically effective.
-A.2 Overview of 100 Secondary Subdomains
+B.2 Overview of 100 Secondary Subdomains
 See Table 8, Table 9 and Table 10.
-Appendix B Prompts
-B.1 Initial Query Generation Prompt
-Initial Query Generation Prompt Generate 10 different writing requests under {domain2} within the context of {domain1}.
-Ensure the requests are as detailed and specific as possible, and reflect realistic user tone and needs.
-Please return in the following JSON format, and do not include anything outside of JSON:
-   "Writing request 1",
-   "Writing request 2",
-Table 6: Initial query generation prompt introduced in Section 3.1.1.
+Table 8: Subdomains in Academic & Engineering and Finance & Business.
 SubDomain Description
@@
 Risk Management Identification, assessment, and prioritization of risks with mitigation strategies
-Sales Report Summary of sales activities, performance, and revenue figures over a given period
-Pitch Deck Visual presentation designed to communicate business ideas or proposals to investors
-Contract Legally binding agreement outlining the terms and conditions for business transactions
+Sales Report Summary of sales activities, performance, and revenue over a specific period
+Pitch Deck Visual presentation for communicating business ideas or proposals to investors
+Contract Legally binding agreement detailing terms and conditions for business transactions
 Tender Document Formal proposal request containing project specifications and bidding instructions
@@
 Requirements Specification Documentation detailing functional and non-functional requirements for a project
-Bid Proposal Formal offer to provide goods or services at a specified price, addressing client needs
-Table 7: Subdomains in Academic & Engineering and Finance & Business.
-Domain Description
+Bid Proposal Formal offer to supply goods or services at a set price, meeting client needs
+Table 9: Subdomains in Politics & Law and Literature & Art.
+Subdomain Description
 Politics & Law
 Legal Opinion Authoritative assessment and guidance on legal matters or questions
-Government Speech Formal address delivered by government officials outlining policies or positions
+Government Speech Formal address by government officials outlining policies or positions
 Judgment Document Official written decision or order issued by a court
@@
 Policy Advocacy Efforts to influence or promote specific policy changes or implementations
-Work Report Detailed account of activities, achievements, and challenges within a specific period
+Work Report Report of activities, achievements, and challenges within a specific period
 Deed Achievement Record highlighting significant accomplishments and contributions
@@
 Book Review Critical evaluation and summary of a book’s content and impact
-Game Design Creation of game mechanics, stories, and interfaces for interactive entertainment
+Game Design Creation of mechanics, stories, and interfaces for games
 Lyric Writing Crafting of words for songs with rhyme and meter considerations
@@
 Poetry Artistic composition using rhythmic and metaphorical language
-Fan Fiction Amateur stories written by enthusiasts featuring characters from existing media
-Table 8: Subdomains in Politics & Law and Literature & Art.
-Domain Description
+Fan Fiction Amateur stories by enthusiasts featuring characters from existing media
+Table 10: Subdomains in Education and Advertising & Marketing.
+SubDomain Description
 Education
@@
 Parent-Teacher Meeting Formal discussion between educators and parents about student progress
-Lesson Plan Structured outline of educational objectives and teaching methods for a class
+Lesson Plan Structured outline of educational objectives and teaching methods
 Teaching Materials Resources used to aid in presenting information to students
@@
 Educational Consulting Professional guidance on educational strategies and systems
-Admissions Promotion Strategies and activities aimed at encouraging enrollment in educational institutions
+Admissions Promotion Strategies and activities to encourage enrollment in educational institutions
 Advertising & Marketing
@@
 Slogans Catchy and memorable phrases designed to convey brand identity
-Table 9: Subdomains in Education and Advertising & Marketing.
-B.2 Guidance Pool
+Appendix C Prompts
+C.1 Query Classification Prompt
+Introduced in Appendix A.1.
+Classification System Prompt You are an expert in query analysis.
+During each iteration, Put current secondary domain tags under ** Domains **. For example, "1. Academic & Engineering" represents a primary domain, while "1a. Thesis Outline" represents a subdomain.
+Query Classification Prompt Please determine which of the following domains the query belongs to and identify any stylistic, formatting, or length requirements.
+**Example of Format Requirement**: Mimicking the format of uploaded documents, adhering to a given outline format, conforming to academic paper formatting standards, etc.
+**Example of Style Requirement**: Suitable for children’s reading, rigorous language, humorous tone, etc.
+**Example of Length Requirement**: Word count, duration, or other constraints related to output size.
+** Query **
+{query}
+** Domains **
+1. Academic & Engineering
+ 1a. Thesis Outline
+ 1b. Literature Review
+ 1r. Others related to Academic & Engineering
+2. Finance & Business
+ 2a. Market Research
+ 2b. Sales Report
+ 2l. Others related to Finance & Business
+7. Other
+ 7a. Others
+** Output format **
+Return in JSON, strictly output according to the following format, do not output other content
+ "domain": ["$x$.yyy",…] // Domains involved, such as "6c. Marketing Letter", "8d. Educational Consulting", if the data is invalid and cannot be determined, return an empty list.
+ ""style": "", // style requirement if present (e.g. "academic format"), else empty string
+ ""format": "", // format specification if present (e.g. "child-friendly tone"), else empty string
+ ""length": "" // length constraint (e.g. "500 words") or empty string
+C.2 Initial Query Generation Prompt
+Introduced in Section 3.1.1. You can specify the language of the generated query in the prompt.
+Query Classification Prompt Generate {NUM} different writing requests under {subdomain} within the context of {primary_domain} in English / Chinese.
+Ensure the requests are as detailed and specific as possible, and reflect realistic user tone and needs.
+Please return in the following JSON format, and do not include anything outside of JSON:
+ "Writing request 1",
+ "Writing request 2",
+C.3 Guidance Pool
+Introduced in Section 3.1.1. Randomly select 0-6 items each time.
 Query Refinement Guidance Pool [
@@
   "Express concisely in one sentence."
-Table 10: Query refinement guidance pool introduced in Section B.5.
-B.3 Query Refine Prompt
+C.4 Query Refine Prompt
+Introduced in Section 3.1.1.
 Query Refinement Prompt Please refine and enhance the original writing requirements in the context of generating content in {domain2} from {domain1} based on the provided guidance.
@@
   "material": "Whether additional reference materials are needed to supplement the writing requirements. If needed, provide suggestions for the materials; if not needed, return"
-Table 11: Query refinement prompt introduced in Section 3.1.1.
-B.4 Criteria Generation Prompt
-System Evaluation Prompt You are an expert evaluator with extensive experience in evaluating the response of a given query.
-Table 12: Evaluation system prompt introduced in Section 3.2.
+C.5 Criteria Generation Prompt
+Introduced in Section 3.2.
+Evaluation System Prompt You are an expert evaluator with extensive experience in evaluating the response of a given query.
 Criteria Generation Prompt Please generate five strict evaluation criteria for assessing the response given the following query.
@@
    "criteria_description": "Description for the first criteria, emphasizing detailed and critical assessment.",
-    "1-2": "Low score description: Clearly deficient in this aspect, with significant issues.",
-    "3-4": "Below average score description: Lacking in several important areas, with noticeable problems.",
-    "5-6": "Average score description: Adequate but not exemplary, meets basic expectations with some minor issues.",
-    "7-8": "Above average score description: Generally strong but with minor shortcomings.",
-    "9-10": "High score description: Outstanding in this aspect, with no noticeable issues." },
-Table 13: Criteria generation prompt introduced in Section 3.2.
-B.5 Rubric-based Scoring Prompt
-Scoring Prompt Evaluate the Response based on the Query and criteria provided.
+   "1-2": "Low score description: Critical deficiencies and major issues that prevent adequate functionality.",
+   "3-4": "Below average score description: Lacking with noticeable shortcomings that impact overall effectiveness and require improvement.",
+   "5-6": "Average score description: Adequate but not exemplary, Baseline performance that meets essential requirements. Most models may achieve this score.",
+   "7-8": "Above average score description: Strong performance characterized by competent execution, though minor refinements are needed to achieve excellence.",
+   "9-10": "High score description: Exceptional performance with all aspects optimally addressed, demonstrating superior effectiveness and quality without any flaws."
+C.6 Rubric-based Scoring Prompt
+Introduced in Section 3.2.
+Evaluation System Prompt You are an expert evaluator with extensive experience in evaluating the response of a given query.
+Since the query and response may be very long, the {criteria} will appear twice: once before and once after the query and response.
+Scoring Prompt Evaluate the Response based on the Query and Criteria provided following the Scoring Rules.
+** Scoring Rules **
+"1-2": "Low score description: Critical deficiencies and major issues that prevent adequate functionality."
+"3-4": "Below average score description: Lacking with noticeable shortcomings that impact overall effectiveness and require improvement."
+"5-6": "Average score description: Adequate but not exemplary, Baseline performance that meets essential requirements. Most models may achieve this score."
+"7-8": "Above average score description: Strong performance characterized by competent execution, though minor refinements are needed to achieve excellence."
+"9-10": "High score description: Exceptional performance with all aspects optimally addressed, demonstrating superior effectiveness and quality without any flaws."
+-Provide reasons for each score by indicating specific strengths or deficiencies within the Response.
+Reference exact text passages to justify the score, ensuring that each reason is concrete and aligns with the criteria requirements while highlighting key gaps from the ideal answer.
+-Be very STRICT and do not be misled by format or length; ensure that the Response is thoroughly evaluated beyond superficial appearances.
+-Carefully discern whether the content of the Response is an illusion, appearing substantial but actually entirely fabricated.
+-Sometimes the model may only provide an introduction or an overview without truly completing the query, which should be considered a failed response. Carefully discern this.
+-Scoring Range: Assign an integer score between 1 to 10
+** Output format **
+(Remove symbols that interfere with JSON parsing, don’t use " inside reason)
+Return the results in the following JSON format, Only output the following JSON format and nothing else:
+  "score": an integer score between 1 to 10,
+  "reason": "Specific and detailed justification for the score using text elements."
 ** Criteria **
@@
 {response}
-Provide your evaluation based on the criteria:
+Provide your evaluation based on the criteria restated below:
 {criteria}
-Provide reasons for each score, indicating where and why any strengths or deficiencies occur within the Response. Reference specific passages or elements from the text to support your justification.
-Ensure that each reason is concrete, with explicit references to the text that align with the criteria requirements.
-Scoring Range: Assign an integer score between 1 to 10
-Output format:
-Return the results in the following JSON format, Only output this JSON format and nothing else:
+** Output format **
+(Remove symbols that interfere with JSON parsing, don’t use " inside reason)
+Return the results in the following JSON format, Only output the following JSON format and nothing else:
   "score": an integer score between 1 to 10,
   "reason": "Specific and detailed justification for the score using text elements."
-Table 14: Scoring prompt introduced in Section 3.2.
-Appendix C Experiment Results
-C.1 Ablation of CoT in Creative content
-To validate the impact of CoT reasoning on creative content generation, we conduct knowledge distillation experiments using DeepSeek-R1.
-Two variants of Qwen-2.5-32B-Instruct are developed: 1) Base model trained with CoT-formatted instructions, 2) ablated version (-w/o CoT) using direct response generation.
-These models are evaluated on the Literature & Art (D5) subset of WritingBench and EQBench, a specialized benchmark for fiction evaluation, to thoroughly examine their capabilities in generating creative content.
-Models WritingBench-D4 EQBench
-Deepseek-R1 8.55 84.99
-Qwen-2.5-32B-Instruct 7.34 48.17
-Qwen-2.5-32B-CoT 8.66 82.48
--w/o CoT 8.49 79.43
-Table 15: Ablation of CoT in creative content on two benchmarks.
-As shown in Table 15, the CoT-enhanced model achieves 8.66 (vs. 8.49) on WritingBench-D5 and 82.48 (vs. 79.43) on EQBench, showing CoT’s effectiveness in generating creative content.
-C.2 Case Study
-Query Create advertising slogans for a pet grooming salon, including: - Main slogan (10-15 words) - Sub-slogan (15-20 words) - Service tagline (8-12 words) - Brand promise (10-15 words) The slogans should reflect professional pet grooming and care services while emphasizing thoughtful care for pet owners.
-The slogans should be catchy, memorable, with warm and friendly wording, and a light, cheerful tone.
-Response **Main Slogan:** Pamper Your Pet, Delight Your Heart.
-*Sub-Slogan:** Expert grooming and loving care, making your furry friend feel fabulous, inside and out.
-*Service Tagline:** Gentle paws, happy pets, stress-free smiles.
-*Brand Promise:** Quality grooming, compassionate care, every time.
-Table 16: Bad case of length requirement. Query from WritingBench with index 539, response generated by Gemini-1.5-Pro.
-Figure 6: Scores of different models across various input lengths on the WritingBench.
-Figure 7: Scores of different models across various output lengths on the WritingBench.
+Appendix D Limitations
+This work faces three primary limitations that warrant consideration.
+First, both the writing model and critic model are primarily trained using conventional SFT approaches, omitting systematic exploration of enhanced optimization strategies.
+While we demonstrate the partial efficacy of the CoT mechanisms in creative domains, their potential remains unexplored compared to established successes in mathematical reasoning tasks.
+Second, our evaluation framework exhibits diminished precision in handling complex multi-dimensional length requirements, including temporal sequencing constraints and section-specific word counts.
+This limitation underscores the necessity for enhanced scoring methodologies that integrate learned metrics with structured rule-based evaluations to better regulate output specifications.
+Third, inherent challenges persist in obtaining reliable pairwise preference annotations for compositional tasks.
+Despite rigorous annotation protocols, human evaluators inevitably introduce subjective biases when assessing two fair-well responses, particularly regarding narrative preferences and contextual interpretations.
+While our consensus-building procedures mitigate some variability, absolute alignment with diverse user preferences remains theoretically unattainable.
+Appendix E Impact
+The introduction of WritingBench and its associated evaluation framework presents significant impacts across multiple dimensions of LLM research and application development:
+  • Comprehensive AI Writing Evaluation Ecosystem WritingBench’s extensive query set spans a wide range of domains and requirements, making it an ideal resource for evaluating both general-purpose writing skills and domain-specific expertise.
+Researchers can use the benchmark to assess a model’s overall versatility or focus on specific domains, such as storytelling, scientific writing, or business communication.
+This dual capability allows for nuanced evaluations that cater to both academic research and practical applications, bridging the gap between theoretical advancements and real-world needs.
+By publicly releasing WritingBench, including its evaluation protocols, criteria generation tools, and critic models, we contribute to greater transparency and reproducibility in the field of LLM research.
+Researchers can replicate experiments, validate findings, and build upon our work to further refine evaluation methodologies.
+Please note that although every released query is manually reviewed to remove harmful content, the responses generated via API calls depend on the underlying model’s behavior, cannot be guaranteed to be fully safe or accurate, and should not be used for commercial purposes.
+  • Facilitating Domain-Specific Research and Development With its inclusion of 100 subdomains, WritingBench provides a unique opportunity for domain-specific optimization.
+For example, industries such as legal writing, medical documentation, or marketing can evaluate model performance on relevant subsets to select the most suitable models for their needs.
+Furthermore, WritingBench’s query construction strategy can be leveraged to generate diverse queries when specialized evaluation datasets are lacking in a particular field.
+This capability supports targeted research the development of highly specialized writing tools.
+We welcome researchers and experts from all fields to join the discussion and collaborate on the development of this benchmark.
+  • Advancing Reinforcement Learning and Adaptive Evaluation Frameworks The query-dependent evaluation framework introduced in WritingBench opens new avenues for integrating adaptive evaluation methods into reinforcement learning pipelines.
+The ability to generate instance-specific criteria dynamically allows for precise scoring mechanisms that can be used to train models through reward-based optimization.
+Additionally, the framework’s capacity to produce pairwise preference data by comparing predicts against adaptive criteria for improving model alignment with human preferences.
+This approach enhances the quality of generated text and accelerates the development of models capable of handling complex, multi-dimensional writing tasks.